### v2

# Dualities of Gaudin models with irregular singularities for general linear Lie (super)algebras

Abstract.
We prove an equivalence between the actions of the Gaudin algebras with irregular singularities for ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ on the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. This establishes a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models. As an application, we show that the Gaudin algebra with irregular singularities for ${\mathfrak{gl}}_{p+m|q+n}$ acts cyclically on each weight space of a certain class of infinite-dimensional modules over a direct sum of Takiff superalgebras over ${\mathfrak{gl}}_{p+m|q+n}$ and that the action is diagonalizable with a simple spectrum under a generic condition. We also study the classical versions of Gaudin algebras with irregular singularities and demonstrate a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for classical Gaudin models.
1. Introduction
Rybnikov [Ry] and Feigin, Frenkel, and Toledano Laredo [FFTL] introduce Gaudin models with irregular singularities associated to any simple Lie algebra over $\mathbb{C}$} based on the works [FF, FFR, Fr1]. Their constructions can also be applied to any general linear Lie (super)algebra.
A number of dualities relating the Gaudin models for a pair of general linear Lie (super)algebras are demonstrated under some restrictions on singularities; see [ChL1, HM, MTV, TU]. Remarkably, Vicedo and Young [VY] establish a duality of Gaudin models for any pair of general linear Lie algebras without imposing such restrictions.
The main goal of this paper is to generalize Vicedo–Young’s duality to the super setting. We will prove a duality of Gaudin models with irregular singularities for the pair $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, where ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ are the general linear Lie (super)algebras defined in Section 2.1.
[…]
${\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}}):=\bigoplus_{i=1}^{{\boldsymbol{\ell}}}{\mathfrak{gl}}_{d}[t_{z_{i}}]/t_{z_{i}}^{\gamma_{i}}{\mathfrak{gl}}_{d}[t_{z_{i}}]$
and
${\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}}):=\bigoplus_{i=1}^{{d^{\prime}}}{\mathfrak{gl}}_{p+m|q+n}[t_{w_{i}}]/t_{w_{i}}^{\xi_{i}}{\mathfrak{gl}}_{p+m|q+n}[t_{w_{i}}].$
Here, for instance, ${\mathfrak{gl}}_{d}[t_{z_{i}}]:={\mathfrak{gl}}_{d}\otimes{\mathbb{C}}[t_{z_{i}}]$ and $t_{z_{i}}^{\gamma_{i}}{\mathfrak{gl}}_{d}[t_{z_{i}}]:={\mathfrak{gl}}_{d}\otimes t_{z_{i}}^{\gamma_{i}}{\mathbb{C}}[t_{z_{i}}]$. The quotient Lie (super)algebras ${\mathfrak{gl}}_{d}[t_{z_{i}}]/t_{z_{i}}^{\gamma_{i}}{\mathfrak{gl}}_{d}[t_{z_{i}}]$ and ${\mathfrak{gl}}_{p+m|q+n}[t_{w_{i}}]/t_{w_{i}}^{\xi_{i}}{\mathfrak{gl}}_{p+m|q+n}[t_{w_{i}}]$ are called Takiff (super)algebras over ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$, respectively; see Section 3.1.
Let $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ denote the Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,{\boldsymbol{\ell}}$, relative to $\bf w$} and $\boldsymbol{\xi}$}, and let $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ denote the Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with singularities of orders $\xi_{i}$ at $w_{i}$, $i=1,\ldots,{d^{\prime}}$, relative to $\bf z$} and $\boldsymbol{\gamma}$}; see Section 3.4, Section 3.5 and (4.15). The algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are commutative subalgebras of the universal enveloping algebras $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ and $U\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$, respectively. Also, for $1\leq i\leq{\boldsymbol{\ell}}$, the singularity at $z_{i}$ is said to be regular if ${\boldsymbol{\gamma}}_{i}=1$ and irregular otherwise. The singularity type (either regular or irregular) of each $w_{j}$ is defined similarly.
Let $\mathcal{F}$} be the polynomial superalgebra generated by $x^{a}_{i}$ and $y^{a}_{r}$, for $i=1,\ldots,m+n$, $r=1,\ldots,p+q$ and $a=1,\ldots,d$. Here the variable $x^{a}_{i}$ (resp., $y^{a}_{r}$) is even for $1\leq i\leq m$ (resp., $1\leq r\leq p$) and is odd otherwise. There are commuting actions of ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ on $\mathcal{F}$} that form a Howe dual pair [CLZ]. The superalgebra $\mathcal{F}$} can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf. [CL, LZ]). Let $\mathcal{D}$ be the corresponding Weyl superalgebra, which is the superalgebra generated by $x^{a}_{i}$ and $y^{a}_{r}$ as well as their derivatives $\frac{\partial}{\partial x^{a}_{i}}$ and $\frac{\partial}{\partial y^{a}_{r}}$. The superalgebra $\mathcal{D}$ naturally acts on $\mathcal{F}$} and is a subalgebra of the endomorphism algebra ${\rm End}({\mathcal{F}})$ of $\mathcal{F}$}. We construct superalgebra homomorphisms $\phi:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)\longrightarrow\mathcal{D}$ and $\varphi:U\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)\longrightarrow\mathcal{D}$ (see Proposition 4.1 and Proposition 4.2) and obtain the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities.
Theorem 1.1 (Theorem 4.5).
$\phi\big(\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\varphi\big(\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big).$
The maps $\phi$ and $\varphi$ induce an action of $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and an action of $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on the Fock space $\mathcal{F}$}, respectively. Theorem 1.1 also says that these actions are equivalent.
A few remarks are in order. In [ChL1] (and also in [HM, MTV, TU] if some of $p,q,m,n$ are set to 0), a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ in the same spirit is established when the entries of $\boldsymbol{\gamma}$} and $\boldsymbol{\xi}$} are all 1 (i.e., the singularities at $z_{i}$’s and $w_{j}$’s are all regular). These conditions mean that the duality considered there concerns only the Gaudin models with the simplest possible singularities. Theorem 1.1 is, however, about a duality of Gaudin models with general singularity types (i.e., no restrictions are imposed on the singularities at $z_{i}$’s and $w_{j}$’s). After specializing to $p=q=n=0$, the theorem recovers the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Vicedo and Young [VY, Theorem 4.8] in the bosonic setting. Furthermore, Theorem 1.1 yields a duality in the fermionic setting by taking $p=q=m=0$; see Corollary 4.7. Specializing to $q=n=0$, the theorem also deduces the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m+p})$ for the bosonic oscillators in which the Fock space $\mathcal{F}$} decomposes into a direct sum of tensor products of infinite-dimensional modules over the general linear Lie algebra ${\mathfrak{gl}}_{m+p}$; see Corollary 4.8.
We can apply Theorem 1.1 to study the Gaudin algebras in the case of ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, where ${\boldsymbol{\ell}}=p+q+m+n$. Using the properties of the action of $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))$ on finite-dimensional ${\mathfrak{gl}}_{d}$-modules, we can obtain some consequences for the action of $\mathcal{A}_{p+m|q+n}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$ on a certain class of ${\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})$-modules. Let $\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{{d^{\prime}}}$, where $M_{i}$ is a certain infinite-dimensional ${\mathfrak{gl}}_{p+m|q+n}[t_{w_{i}}]/t_{w_{i}}^{\xi_{i}}{\mathfrak{gl}}_{p+m|q+n}[t_{w_{i}}]$-module for $1\leq i\leq{d^{\prime}}$. We will prove in Theorem 4.13 that every weight space of $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a cyclic $\mathcal{A}_{p+m|q+n}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$-module, and that $\mathcal{A}_{p+m|q+n}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$ is diagonalizable with a simple spectrum on the weight space for generic $\bf w$} and $\bf z$}.
In this paper, we also study the classical versions of Gaudin models. Let $\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ and $\mathcal{S}\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$ be the supersymmetric algebras of ${\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})$ and ${\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})$, respectively. They are also Poisson superalgebras and may be viewed as classical counterparts of $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ and $U\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$, respectively. Let $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\subseteq\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ be the classical Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,{\boldsymbol{\ell}}$, relative to $\bf w$} and $\boldsymbol{\xi}$}, and let $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\subseteq\mathcal{S}\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)$ be the classical Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with singularities of orders $\xi_{i}$ at $w_{i}$, $i=1,\ldots,{d^{\prime}}$, relative to $\bf z$} and $\boldsymbol{\gamma}$}; see Section 5.2 and Section 5.3. The classical Gaudin algebras $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are Poisson-commutative.
Let $\overline{\mathcal{D}}$ be the polynomial superalgebra generated by the variables $x^{a}_{i}$, $y^{a}_{r}$, $p_{x^{a}_{i}}$ and $p_{y^{a}_{r}}$, for $i=1,\ldots,m+n$, $r=1,\ldots,p+q$ and $a=1,\ldots,d$, where $x^{a}_{i}$ and $p_{x^{a}_{i}}$ (resp., $y^{a}_{r}$ and $p_{y^{a}_{r}}$) are even for $1\leq i\leq m$ (resp., $1\leq r\leq p$) and are odd otherwise. It is a Poisson superalgebra with the Poisson bracket defined by (5.8) and (5.9). There also exist Poisson superalgebra homomorphisms $\overline{\phi}:\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)\longrightarrow\overline{\mathcal{D}}$ and $\overline{\varphi}:\mathcal{S}\big({\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})\big)\longrightarrow\overline{\mathcal{D}}$; see Proposition 5.2. The duality in Theorem 1.1 has a classical version, which we call the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for classical Gaudin models with irregular singularities.
Theorem 1.2 (Theorem 5.3).
$\overline{\phi}\big(\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\overline{\varphi}\big(\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big)$.
Theorem 1.2 recovers the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ for classical Gaudin models due to Vicedo and Young [VY, Theorem 3.2] by taking $p=q=n=0$ and the fermionic counterpart [VY, Theorem 3.4] by taking $p=q=m=0$.
We organize the paper as follows. In Section 2, we review the background materials needed in this paper. In Section 3, we discuss Gaudin models with irregular singularities and their fundamental properties. In Section 4, we introduce a joint action of the Gaudin algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on the Fock space $\mathcal{F}$}. We prove Theorem 1.1, establishing a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models. Specializing to ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, we give an application to a class of modules over ${\mathfrak{gl}}_{p+m|q+n}({{\bf w},{\boldsymbol{\xi}}})$ and obtain Theorem 4.13. In Section 5, we introduce the classical Gaudin algebras $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\overline{\mathcal{A}}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$. We prove Theorem 1.2 and obtain a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for classical Gaudin models.
Notations. Throughout the paper, the symbol $\mathbb{Z}$} (resp., $\mathbb{N}$} and $\mathbb{Z}$$_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol $\mathbb{C}$} for the field of complex numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2. All vector spaces, algebras, tensor products, etc., are over $\mathbb{C}$}. We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
2. Preliminaries
In this section, we review general linear Lie (super)algebras, column determinants, Berezinians, and pseudo-differential operators.
2.1. The general linear Lie (super)algebra
For $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero, let
$\mathbb{I}=\{i\in{\mathbb{N}}\,|\,1\leq i\leq p+q+m+n\}.$
For $i\in\mathbb{I}$, define $|i|\in{\mathbb{Z}}_{2}$ by
$|i|=\begin{cases}\,\bar{0}&\ \ \mbox{if}\ \ i\in\{1,\ldots,p\}\cup\{p+q+1,\ldots,p+q+m\};\\ \,\bar{1}&\ \ \mbox{otherwise}.\end{cases}$
Let $\{e_{i}\,|\,i\in\mathbb{I}\}$ be a basis for the superspace ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ such that $\{e_{i}\,|\,1\leq i\leq p+q\}$ and $\{e_{p+q+i}\,|\,1\leq i\leq m+n\}$ are respectively the standard homogeneous bases for ${\mathbb{C}}^{p|q}$ and ${\mathbb{C}}^{m|n}$. In other words, the parity of $e_{i}$ is given by $|e_{i}|=|i|$ for $i\in\mathbb{I}$.
For any $i,j\in\mathbb{I}$, let $E^{i}_{j}$ denote the $\mathbb{C}$}-linear endomorphism on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ defined by
$E^{i}_{j}(e_{k})=\delta_{j,k}e_{i}\quad\mbox{ for $k\in\mathbb{I}$,}$
where $\delta$ is the Kronecker delta. The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$. The superspace of $\mathbb{C}$}-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ is a Lie superalgebra, called a general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$, with commutation relations given by
$E^{i}_{j},E^{k}_{l}]=\delta_{j,k}E^{i}_{l}-(-1)^{(|i|+|j|)(|k|+|l|)}\delta_{i,l}E^{k}_{j}\qquad\mbox{for $i,j,k,l\in\mathbb{I}$.}$
Note that $\{E^{i}_{j}\,|\,i,j\in\mathbb{I}\}$ is a homogeneous basis for ${\mathfrak{gl}}_{p+m|q+n}$.
For the rest of the paper, we use the symbol $\diamondsuit:={p+m|q+n}$. Write
${\mathfrak{gl}}_{\diamondsuit}={\mathfrak{gl}}_{p+m|q+n}.$
Let $\mathfrak{b}_{\diamondsuit}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of $\mathfrak{gl}$$_{\diamondsuit}$. The corresponding Cartan subalgebra $\mathfrak{h}$$_{\diamondsuit}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$, and the dual basis in $\mathfrak{h}$$_{\diamondsuit}^{*}$ is denoted by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$, where the parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$. Note that the Borel subalgebra $\mathfrak{b}_{\diamondsuit}$ is not the standard one unless some of $p,q,m,n$ are set to 0.
For $m=d$ and $p=q=n=0$, we write
${\mathfrak{gl}}_{d}={\mathfrak{gl}}_{d|0}$
and
$\hskip 28.45274pte_{ij}=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$
2.2. Column determinants and Berezinians
Let $\mathcal{A}$ be an associative unital superalgebra over $\mathbb{C}$}. The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$. The superalgebra $\mathcal{A}$ is naturally a Lie superalgebra with supercommutator
(2.1) $a,b]:=ab-(-1)^{|a||b|}ba$
for homogeneous elements $a,b\in\mathcal{A}$.
Fix $k\in{\mathbb{N}}$. For any $k\times k$ matrix $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
${\rm cdet}\,(A)=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$, and $l(\sigma)$ denotes the length of $\sigma$. The row determinant of $A$ is defined to be
${\rm rdet}\,(A)=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{1,\sigma(1)}\ldots a_{k,\sigma(k)}.$
Evidently, ${\rm rdet}\,(A)={\rm cdet}\,(A^{t})$, where $A^{t}$ is the transpose of $A$. If all the entries of $A$ commute, then the column determinant and the row determinant of $A$ coincide, and we define the determinant of $A$ to be
$\det(A)={\rm cdet}\,(A)={\rm rdet}\,(A).$
Assume that $A$ has a two-sided inverse $A^{-1}=\big[\widetilde{a}_{i,j}\big]$. For $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$. If $k\geq 2$, then
$|A|_{ij}=a_{i,j}-r_{i}^{j}(A^{ij})^{-1}c_{j}^{i},$
where $A^{ij}$ is the $(k-1)\times(k-1)$ submatrix of $A$ obtained by deleting the $i$th row and the $j$th column of $A$ and is assumed to be invertible, $r_{i}^{j}$ is the $i$th row of $A$ with $a_{i,j}$ removed, and $c_{j}^{i}$ is the $j$th column of $A$ with $a_{i,j}$ removed (see [GGRW, Proposition 1.2.6]). Following [GGRW], it is convenient to write
$|A|_{ij}=\begin{vmatrix}a_{1,1}&\ldots&a_{1,j}&\ldots&a_{1,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,j}$}&\ldots&a_{i,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{k,1}&\ldots&a_{k,j}&\ldots&a_{k,k}\end{vmatrix}$.
For $i=1,\ldots,k$, we define
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix}$,\
called the principal quasiminors of $A$.
Let $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ be a $k\times k$ matrix over $\mathcal{A}$. For any nonempty subset $P=\{i_{1}<\ldots<i_{\ell}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big[a_{i,j}\big]_{i,j\in P}$ is called a standard submatrix of $A$. Moreover, we say that $A$ is sufficiently invertible if every principal quasiminor of $A$ is well defined, and that $A$ is amply invertible if each of its standard submatrices is sufficiently invertible.
Let $(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1. We call such a sequence a $0^{m}1^{n}$-sequence. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 0’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
Let ${\bf s}=(s_{1},\ldots,s_{m+n})\in\mathcal{S}_{{m|n}}$. For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A:=[a_{i,j}]$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big[a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big]$. We say that $A$ is of type $\bf s$} if $a_{i,j}$ is a homogeneous element of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for $1\leq i,j\leq m+n$.
Following the definition of [HM], for any $(m+n)\times(m+n)$ sufficiently invertible matrix $A$ of type $\bf s$} over $\mathcal{A}$, the Berezinian of type $\bf s$} of $A$ is defined to be
${\rm Ber}^{\bf s}\,(A)=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $1\leq i\leq m+n$. The original formulation of Berezinians, in the case where $\mathcal{A}$ is supercommutative, was given in [Ber].
Proposition 2.1 (cf. [HM, Proposition 3.5]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible matrix of type $\bf s$} over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
(2.2) $A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix}$,
where $W,X,Y,Z$ are respectively $k\times k$, $k\times(m+n-k)$, $(m+n-k)\times k$, and $(m+n-k)\times(m+n-k)$ matrices. Then $W$ and $Z-YW^{-1}X$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,(A)={\rm Ber}^{{\bf s}^{\prime}}(W)\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big(Z-YW^{-1}X\big).$
An $(m+n)\times(m+n)$ matrix $A=\big[a_{i,j}\big]$ over $\mathcal{A}$ is called a Manin matrix of type $\bf s$} if $A$ is a matrix of type $\bf s$} satisfying
$a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]$
for $1\leq i,j,k,l\leq m+n$. Evidently, if $\mathcal{A}$ is supercommutative, then any $(m+n)\times(m+n)$ matrix of type $\bf s$} is automatically a Manin matrix.
Let us recall some useful facts about Manin matrices.
Proposition 2.2 (cf. [HM, Section 3]).
Let $A$ be an $(m+n)\times(m+n)$ Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
  1. (i)
    If $P=\{i_{1}<\ldots<i_{\ell}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ of $A$ is a Manin matrix of type ${{\bf s}}_{P}:=(s_{i_{1}},\ldots s_{i_{\ell}})$.
  2. (ii)
    For any $\sigma\in\mathfrak{S}_{m+n}$, $A^{\sigma}$ is a Manin matrix of type ${\bf s}^{\sigma}$.
Proposition 2.3 ([HM, Proposition 3.6]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Then ${\rm Ber}^{{\bf s}^{\sigma}}(A^{\sigma})={\rm Ber}^{\bf s}\,(A)$ for any $\sigma\in\mathfrak{S}_{m+n}$.
Proposition 2.4 ([ChL2, Proposition 4.4]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Write $A$ as in (2.2). Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,(A)={\rm Ber}^{{\bf s}^{\prime\prime}}(Z)\cdot{\rm Ber}^{{\bf s}^{\prime}}\big(W-XZ^{-1}Y\big).$
Proposition 2.5 ([CFR, Lemma 8]).
Let ${\bf s}_{0}=(0^{m})$, and let $A$ be an $m\times m$ sufficiently invertible Manin matrix of type ${\bf s}_{0}$ over $\mathcal{A}$. Then ${\rm Ber}\,^{{\bf s}_{0}}(A)={\rm cdet}\,(A)$.
Proposition 2.6.
Suppose that $\mathcal{A}$ is supercommutative. Let ${\bf s}_{1}=(1^{n})$, and let $A$ be an $n\times n$ sufficiently invertible matrix of type ${\bf s}_{1}$ over $\mathcal{A}$. Then ${\rm Ber}\,^{{\bf s}_{1}}(A)=[\det(A)]^{-1}$.
Proof.
By hypothesis, $A$ is a matrix of type $(0^{n})$ with commuting entries. By Proposition 2.5, $\det(A)=d_{1}(A)\ldots d_{n}(A)$. It follows that $\det(A)]^{-1}=d_{1}(A)^{-1}\ldots d_{n}(A)^{-1}={\rm Ber}\,^{{\bf s}_{1}}(A)$ as the elements $d_{i}(A)$ commute. ∎
2.3. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra over $\mathbb{C}$}, and let $z$ be an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ (resp., $\mathcal{A}(\!(z)\!)$) the superalgebra of formal power series (resp., formal Laurent series) in $z$ with coefficients in $\mathcal{A}$.
The superalgebra $\mathcal{A}(\!(z^{-1})\!)$ is equipped with an even derivation $\partial_{z}$ defined by
$\partial_{z}\left(\sum_{i=-\infty}^{r}a_{i}z^{i}\right)=\sum_{i=-\infty}^{r}ia_{i}z^{i-1},\qquad r\in{\mathbb{Z}}\quad\text{and}\quad a_{i}\in\mathcal{A}.$
Let $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ be the superspace of formal Laurent series in $\partial_{z}^{-1}$ with coefficients in $\mathcal{A}(\!(z^{-1})\!)$. The multiplication on $\mathcal{A}(\!(z^{-1})\!)$ extends to a multiplication on $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ by the rules
(2.3) $\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{i}\phi=\sum_{k=0}^{\infty}\binom{i}{k}\,\partial_{z}^{k}(\phi)\,\partial_{z}^{i-k},\quad\mbox{for $\phi\in\mathcal{A}(\!(z^{-1})\!)$ and $i\in{\mathbb{Z}}$.}$
Here ${\binom{i}{k}:=\frac{i(i-1)\ldots(i-k+1)}{k!}}$ for $i\in{\mathbb{Z}}$ and $k\in{\mathbb{Z}}_{+}$.
Lemma 2.7.
$\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ is an associative unital superalgebra over $\mathbb{C}$}.
Proof.
We need only show that the associative law for multiplication holds in $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$. To establish the law, it suffices to show that
(2.4) $z^{i_{1}}\partial_{z}^{j_{1}}\big(z^{i_{2}}\partial_{z}^{j_{2}}z^{i_{3}}\partial_{z}^{j_{3}}\big)=\big(z^{i_{1}}\partial_{z}^{j_{1}}z^{i_{2}}\partial_{z}^{j_{2}}\big)z^{i_{3}}\partial_{z}^{j_{3}}$
for $i_{r},j_{r}\in{\mathbb{Z}}$, $r=1,2,3$. By (2.3),
$z^{i_{1}}\partial_{z}^{j_{1}}\big(z^{i_{2}}\partial_{z}^{j_{2}}z^{i_{3}}\partial_{z}^{j_{3}}\big)=\sum_{k=0}^{\infty}a_{k}z^{i_{1}+i_{2}+i_{3}-k}\partial_{z}^{j_{1}+j_{2}+j_{3}-k},$
where
$a_{k}=\sum_{r=0}^{k}\binom{j_{1}}{r}\binom{i_{2}+i_{3}-k+r}{r}\binom{j_{2}}{k-r}\binom{i_{3}}{k-r}r!(k-r)!,$
and
$\big(z^{i_{1}}\partial_{z}^{j_{1}}z^{i_{2}}\partial_{z}^{j_{2}}\big)z^{i_{3}}\partial_{z}^{j_{3}}=\sum_{k=0}^{\infty}b_{k}z^{i_{1}+i_{2}+i_{3}-k}\partial_{z}^{j_{1}+j_{2}+j_{3}-k},$
where
$b_{k}=\sum_{r=0}^{k}\binom{j_{1}}{r}\binom{i_{2}}{r}\binom{j_{1}+j_{2}-r}{k-r}\binom{i_{3}}{k-r}r!(k-r)!.$
Using Vandermonde’s identity:
$\binom{i+j}{r}=\sum_{s=0}^{r}\binom{i}{s}\binom{j}{r-s},\qquad i,j\in{\mathbb{Z}}\,\,\text{ and }\,\,r\in{\mathbb{Z}}_{+},$
we have for $k\in{\mathbb{Z}}_{+}$,
$a_{k}$ $=$ $\sum_{r=0}^{k}\binom{j_{1}}{r}\binom{j_{2}}{k-r}\left(\sum_{s=0}^{r}\binom{i_{2}}{r-s}\binom{i_{3}-k+r}{s}\right)\binom{i_{3}}{k-r}r!(k-r)!$
$=$ $\sum_{r=0}^{k}\sum_{s=0}^{r}\binom{j_{1}}{r}\binom{j_{2}}{k-r}\binom{i_{2}}{r-s}\binom{i_{3}}{k-r+s}\frac{r!(k-r+s)!}{s!}$
and
$b_{k}$ $=$ $\sum_{r=0}^{k}\binom{i_{2}}{r}\binom{i_{3}}{k-r}\binom{j_{1}}{r}\left(\sum_{s=0}^{k-r}\binom{j_{1}-r}{s}\binom{j_{2}}{k-r-s}\right)r!(k-r)!$
$=$ $\sum_{r=0}^{k}\sum_{s=0}^{k-r}\binom{i_{2}}{r}\binom{i_{3}}{k-r}\binom{j_{1}}{r+s}\binom{j_{2}}{k-r-s}\frac{(r+s)!(k-r)!}{s!}$
$=$ $\sum_{\ell=0}^{k}\sum_{s=0}^{\ell}\binom{i_{2}}{\ell-s}\binom{i_{3}}{k-\ell+s}\binom{j_{1}}{\ell}\binom{j_{2}}{k-\ell}\frac{\ell!(k-\ell+s)!}{s!}.$
Thus, $a_{k}=b_{k}$ for $k\in{\mathbb{Z}}_{+}$, and (2.4) is proved. ∎
The elements of the superalgebra $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ are called pseudo-differential operators. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the subspace of $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$ consisting of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}\partial_{z}^{j},\qquad r,s\in{\mathbb{Z}}\quad\text{and}\quad a_{ij}\in\mathcal{A}.$
It is clear that $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is closed under multiplication, and so it is a subalgebra of $\mathcal{A}(\!(z^{-1})\!)(\!(\partial_{z}^{-1})\!)$. Applying the rules (2.3), we see that any expression of the form ${\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}\partial_{z}^{i}z^{j}}$, where $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$, is an element of $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$. In particular,
(2.5) $\sum_{k=0}^{\infty}(-1)^{k}\binom{i}{k}\binom{j}{k}k!\,\partial_{z}^{j-k}\,z^{i-k}=z^{i}\partial_{z}^{j},\qquad\mbox{$i,j\in{\mathbb{Z}}$.}$
Conversely, any element of $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ can be written as ${\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}\partial_{z}^{i}z^{j}}$ for some $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$ due to (2.5).
It is easy to see from (2.3) and (2.5) that the identity map on the superalgebra $\mathcal{A}$ extends to a superalgebra automorphism
(2.6) ${\boldsymbol{\omega}}:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$
such that ${\boldsymbol{\omega}}(z)=\partial_{z}$ and ${\boldsymbol{\omega}}(\partial_{z})=-z$. Also, $\boldsymbol{\omega}$} has order 4.
Let $z$ and $w$ be commuting even variables. For later use, we let $\mathcal{A}(\!(z^{-1},w^{-1})\!)$ denote the superalgebra of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}w^{j},\qquad r,s\in{\mathbb{Z}}\quad\text{and}\quad a_{ij}\in\mathcal{A}.$
3. Gaudin models with irregular singularities
In this section, we recall Takiff superalgebras and introduce evaluation maps of higher orders. The primary purpose is to discuss Gaudin models with irregular singularities [Ry, FFTL] and their fundamental properties.
3.1. Takiff superalgebras
For any Lie superalgebra $\mathfrak{g}$}, we denote by $U({\mathfrak{g}})$ the universal enveloping algebra of $\mathfrak{g}$}. For any even variable $t$, the loop algebra ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ is defined to be the Lie superalgebra with commutation relations
$\hskip 28.45274pt\big[A\otimes t^{r},B\otimes t^{s}\big]=[A,B]\otimes t^{r+s}\qquad\mbox{for $A,B\in{\mathfrak{g}}$ and $r,s\in{\mathbb{Z}}$.}$
Here $[A,B]$ is the supercommutator of $A$ and $B$. The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ is a subalgebra of ${\mathfrak{g}}[t,t^{-1}]$. We identify the Lie superalgebra $\mathfrak{g}$} with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$.
For any $\gamma\in{\mathbb{N}}$, $t^{\gamma}{\mathfrak{g}}[t]:={\mathfrak{g}}\otimes t^{\gamma}{\mathbb{C}}[t]$ is an ideal of ${\mathfrak{g}}[t]$. The corresponding quotient Lie superalgebra ${\mathfrak{g}}^{(\gamma)}[t]:={\mathfrak{g}}[t]/t^{\gamma}{\mathfrak{g}}[t]$ is called a (generalized) Takiff superalgebra over $\mathfrak{g}$}. For the sake of simplicity, we write
$A\otimes\overline{t}^{i}=A\otimes t^{i}+t^{\gamma}{\mathfrak{g}}[t]\qquad\mbox{for $i\in{\mathbb{Z}}_{+}$}.$
3.2. Evaluation maps of higher orders
Let $\mathfrak{g}$} be a Lie (super)algebra and $a\in{\mathbb{C}}$. There is an evaluation homomorphism ${\rm ev}_{a}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})$ given by
${\rm ev}_{a}(A\otimes t^{r})=a^{r}A\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}_{+}$}.$
For $\gamma\in{\mathbb{N}}$, write
$t_{a}=t-a\qquad\text{and}\qquad{\mathfrak{g}}(a,\gamma)={\mathfrak{g}}^{(\gamma)}[t_{a}]$
There is a Lie superalgebra homomorphism ${\mathfrak{g}}[t]\longrightarrow{\mathfrak{g}}(a,\gamma)$, defined by
$A\otimes t^{r}\mapsto\sum_{i=0}^{r}\binom{r}{i}a^{r-i}(A\otimes\overline{t}_{a}^{i}),$
for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}_{+}$. It extends to a superalgebra homomorphism
${\rm ev}_{(a,\gamma)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}(a,\gamma)),$
which we call the evaluation map of order $\gamma$ at $a$. If $\gamma=1$, we have the commutative diagram
$U({\mathfrak{g}}[t])$$U({\mathfrak{g}}(a,1))$$U({\mathfrak{g}})$$\scriptstyle{{\rm ev}_{a}}$$\scriptstyle{{\rm ev}_{(a,1)}}\scriptstyle{\cong}$
Fix $\ell\in{\mathbb{N}}$. For ${\bf z}:=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$ and ${\boldsymbol{\gamma}}:=(\gamma_{1},\ldots,\gamma_{\ell})\in{\mathbb{N}}^{\ell}$, consider the direct sum of Takiff superalgebras
${\mathfrak{g}}({{\bf z},{\boldsymbol{\gamma}}}):=\bigoplus_{i=1}^{\ell}{\mathfrak{g}}(z_{i},\gamma_{i}).$
These $z_{i}$’s correspond to the singularities for the Gaudin algebras to be defined in Section 3.4 and Section 3.5.
Let $\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])^{\otimes\ell}$ be the $(\ell-1)$-fold coproduct on $U({\mathfrak{g}}[t])$. The composite map
$\underline{\rm ev}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}}$
is called the evaluation map at $\bf z$}. On the other hand, we also have the map
$\underline{{\rm ev}}_{({\bf z},{\boldsymbol{\gamma}})}:=\left({\rm ev}_{(z_{1},\gamma_{1})}\otimes\ldots\otimes{\rm ev}_{(z_{\ell},\gamma_{\ell})}\right)\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}({{\bf z},{\boldsymbol{\gamma}}})),$
called the evaluation map of order $\boldsymbol{\gamma}$} at $\bf z$}. If ${\boldsymbol{\gamma}}=(1^{\ell})$, i.e., $\gamma_{i}=1$ for $1\leq i\leq\ell$, then we have the commutative diagram
$U({\mathfrak{g}}[t])$$U\left({\mathfrak{g}}\!\left({\bf z},(1^{\ell})\right)\right)$$U({\mathfrak{g}})^{\otimes\ell}$$\scriptstyle{\underline{\rm ev}_{\bf z}}$$\scriptstyle{\underline{{\rm ev}}_{\left({\bf z},(1^{\ell})\right)}}\scriptstyle{\cong}$
3.3. Feigin–Frenkel centers
Let $t$ be an even variable. For a general linear Lie (super)algebra $\mathfrak{g}$}, let $V_{{\rm crit}}({\mathfrak{g}})$ be the universal affine vertex algebra at the critical level associated to the affine Lie (super)algebra $\widehat{{\mathfrak{g}}}:={\mathfrak{g}}[t,t^{-1}]\oplus{\mathbb{C}}K$; see [Fr2, FBZ, Mo, MR] for details. Note that $V_{{\rm crit}}({\mathfrak{g}})$ is a $\widehat{{\mathfrak{g}}}-$module. The center of the vertex algebra $V_{{\rm crit}}({\mathfrak{g}})$ is given by
${\mathfrak{z}}(\widehat{{\mathfrak{g}}})=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{g}})\;\middle|\;{\mathfrak{g}}[t]v=0\!\,\right\},$
called the Feigin–Frenkel center. By the Poincaré–Birkhoff–Witt theorem, we identify $V_{{\rm crit}}({\mathfrak{g}})$ with $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ (as superspaces). Moreover, the (super)algebra $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is equipped with the (even) derivation ${\rm T}:=-d/dt$ defined by
(3.1) ${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A\otimes t^{-r})=rA\otimes t^{-r-1}$
[…]
Suppose $\mu\in{\mathfrak{g}}^{*}$ vanishes on the odd part of $\mathfrak{g}$}. Then there is a superalgebra homomorphism
$\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
given by
$\hskip 28.45274pt\Psi^{\mu}(A\otimes t^{-r})=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\quad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$
The evaluation map $\underline{{\rm ev}}_{({\bf z},{\boldsymbol{\gamma}})}$ extends to
$\underline{{\rm ev}}_{({\bf z},{\boldsymbol{\gamma}})}:U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]\longrightarrow U({\mathfrak{g}}({{\bf z},{\boldsymbol{\gamma}}}))[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
by setting $\underline{{\rm ev}}_{({\bf z},{\boldsymbol{\gamma}})}(z^{-1})=z^{-1}$. We have the composite map
$\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}:=\underline{{\rm ev}}_{({\bf z},{\boldsymbol{\gamma}})}\circ\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U\big({\mathfrak{g}}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt].$
Using the identity
$(t-z)^{-1}=-\frac{1}{z-z_{i}}\left(1-\frac{t_{z_{i}}}{z-z_{i}}\right)^{-1}=-\sum_{k=0}^{\infty}\frac{t_{z_{i}}^{k}}{(z-z_{i})^{k+1}}$
for $1\leq i\leq\ell$, we find that
(3.2) $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\left(A\otimes t^{-1}\right)=-\sum_{i=1}^{\ell}\sum_{k=0}^{\gamma_{i}-1}\frac{A\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}}+\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}$.}$
Let
$\tau=-\partial_{t},$
where $t$ and $\partial_{t}$ satisfy the rules similar to (2.3). The map $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}$ extends to a superalgebra homomorphism
$\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}:U(t^{-1}{\mathfrak{g}}[t^{-1}])(\!(\tau^{-1})\!)\longrightarrow U\big({\mathfrak{g}}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)$
defined by
$\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\left(\sum_{i=-\infty}^{r}a_{i}\tau^{i}\right)=\sum_{i=-\infty}^{r}\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(a_{i})\partial_{z}^{i}$
for $a_{i}\in U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and $r\in{\mathbb{Z}}$.
3.4. Gaudin algebras for ${\mathfrak{gl}}_{d}$
Let
$\mathcal{T}_{d}=\Big[\delta_{i,j}\tau+e_{ij}\otimes t^{-1}\Big]_{i,j=1,\ldots,d},$
which is a $d\times d$ Manin matrix over $U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]$. We have an expansion
${\rm cdet}\,(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}a_{i}\tau^{i-1}$
for some $a_{1},\ldots,a_{d}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$.
Theorem 3.1 ([CM, Theorem 3.1] (cf. [CT])).
The set $\{{\rm T}^{r}a_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}\}$ is algebraically independent, and ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})={\mathbb{C}}\!\left[{\rm T}^{r}a_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}\,\!\right].$
Fix ${\bf z}:=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$ and ${\boldsymbol{\gamma}}:=(\gamma_{1},\ldots,\gamma_{\ell})\in{\mathbb{N}}^{\ell}$. For any $\mu\in{\mathfrak{gl}}_{d}^{*}$, let
${\mathcal{L}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big(\delta_{i,j}\tau+e_{ij}\otimes t^{-1}\big)\Big]_{i,j=1,\ldots,d}.$
Since $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}$ is a homomorphism, we have
${\rm cdet}\,({\mathcal{L}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}}))=\partial_{z}^{d}+\sum_{i=1}^{d}a_{i}(z)\partial_{z}^{i-1},$
where $a_{i}(z):=\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(a_{i})\in U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt]$.
[…]
Remark 3.2.
The algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ coincides with the subalgebra of $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\in U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt]$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. This follows from Theorem 3.1 and $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big({\rm T}(S)\big)=d/dz\big(\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\big)$ (cf. [ChL1, Proposition 5.9]).
For any $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$-module $V$, let $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ denote the image of the Gaudin algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ in ${\rm End}(V)$. It is called the Gaudin algebra of $V$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$.
As in (2.6), the identity map on the algebra $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ extends to an automorphism
${\boldsymbol{\omega}}:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)$
such that ${\boldsymbol{\omega}}(z)=\partial_{z}$ and ${\boldsymbol{\omega}}(\partial_{z})=-z$. Evidently,
${\mathcal{L}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\delta_{i,j}\partial_{z}+\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big(e_{ij}\otimes t^{-1}\big)\Big]_{i,j=1,\ldots,d}.$
Set
(3.3) $\widehat{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=-\Big[{\boldsymbol{\omega}}\big(\delta_{i,j}\partial_{z}-\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big(e_{ji}\otimes t^{-1}\big)\Big]_{i,j=1,\ldots,d}.$
We have the following (see [VY, the discussion in Section 4.2]).
Proposition 3.3.
The column determinant ${\rm cdet}\,\big(\widehat{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})\big)$ has an expansion
${\rm cdet}\,\big(\widehat{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})\big)=z^{d}+\sum_{i=1}^{d}\widehat{a}_{i}(\partial_{z})z^{i-1}$
for some $\widehat{a}_{i}(\partial_{z})\in U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[\partial_{z}^{-1}]\kern-1.1pt]$. Moreover, the algebra $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ is generated by the coefficients of the series $\widehat{a}_{i}(\partial_{z})$ for $i=1,\ldots,d$.
3.5. Gaudin algebras for ${\mathfrak{gl}}_{p+m|q+n}$
Recall the symbol $\diamondsuit={p+m|q+n}$ and fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{\diamondsuit}$. Let
$\mathcal{T}_{\diamondsuit}=\Big[\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}\otimes t^{-1}\Big]_{i,j\in\mathbb{I}}.$
We know that $\mathcal{T}_{\diamondsuit}$ is an amply invertible $(p+q+m+n)\times(p+q+m+n)$ Manin matrix of type $\bf s$} over $U(t^{-1}{\mathfrak{gl}}_{\diamondsuit}[t^{-1}])(\!(\tau^{-1})\!)$. We have an expansion
${\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{\diamondsuit}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}\tau^{i},$
[…]
$\{{\rm T}^{r}b_{i}\,|\,i\leq p+m-q-n,\,i\in{\mathbb{Z}},\,r\in{\mathbb{Z}}_{+}\}.$
[…]
Fix $\bf z$$\in{\mathbb{C}}^{\ell}$ and $\boldsymbol{\gamma}$$\in{\mathbb{N}}^{\ell}$. For any $\mu\in{\mathfrak{gl}}_{\diamondsuit}^{*}$ which vanishes on the odd part of $\mathfrak{gl}$$_{\diamondsuit}$, let
(3.4) ${\mathcal{L}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big(\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}\otimes t^{-1}\big)\Big]_{i,j\in\mathbb{I}}.$
Then
${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}(z)\partial_{z}^{i},$
which belongs to $U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)$. Here $b_{i}(z):=\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(b_{i})$.
Let $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of the series $b_{i}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. We call $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ the Gaudin algebra for $\mathfrak{gl}$$_{\diamondsuit}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$. Also, for $1\leq i\leq\ell$, the singularity at $z_{i}$ is said to be regular if $\gamma_{i}=1$ and irregular otherwise.
Remark 3.4.
  1. (i)
    The algebra $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ equals the subalgebra of $U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)$ for $S\in\hat{\mathfrak{z}}_{\diamondsuit}$ (as $\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}\big({\rm T}(S)\big)=d/dz\big(\Psi^{\mu}_{({\bf z},{\boldsymbol{\gamma}})}(S)\big)$) (cf. [ChL1, Proposition 5.9]).
  2. (ii)
    We do not know whether Remark 3.2 holds for $\mathfrak{gl}$$_{\diamondsuit}$ in general as it is still a conjecture that $^{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{\diamondsuit})=\hat{\mathfrak{z}}_{\diamondsuit}$. (Note: The conjecture has been shown to be valid for ${\mathfrak{gl}}_{1|1}$ [$M$] and ${\mathfrak{gl}}_{2|1}$ [AN].)
[…]
For any $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$-module $V$, let $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ denote the image of the Gaudin algebra $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ in ${\rm End}(V)$. We call $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})_{V}$ the Gaudin algebra of $V$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$.
4. A duality for Gaudin models
In this section, we construct a joint action of the Gaudin algebras with irregular singularities for ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ on the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. We establish a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models and give an application to the action of the Gaudin algebra for ${\mathfrak{gl}}_{p+m|q+n}$ on a certain class of modules over a direct sum of Takiff superalgebras over ${\mathfrak{gl}}_{p+m|q+n}$.
4.1. Actions on Fock spaces
Let $\mathcal{F}$} be the polynomial superalgebra generated by the variables $x^{a}_{i}$ and $y^{a}_{r}$, for $i=1,\ldots,m+n$, $r=1,\ldots,p+q$ and $a=1,\ldots,d$, where $x^{a}_{i}$ (resp., $y^{a}_{r}$) are even for $1\leq i\leq m$ (resp., $1\leq r\leq p$) and are odd otherwise. There are commuting actions of ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ on $\mathcal{F}$} that form a Howe dual pair [CLZ]. The superalgebra $\mathcal{F}$} may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf. [CL, LZ]). Let $\mathcal{D}$ be the corresponding Weyl superalgebra. That is, $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i}$ and $y^{a}_{r}$ as well as their derivatives
$\partial_{x^{a}_{i}}:=\frac{\partial}{\partial x^{a}_{i}}\qquad\text{and}\qquad\partial_{y^{a}_{r}}:=\frac{\partial}{\partial y^{a}_{r}}$
for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$. The superalgebra $\mathcal{D}$ is naturally a Lie superalgebra with supercommutator $\cdot$$\cdot$] defined as in (2.1). Clearly, $\mathcal{D}$ acts naturally on $\mathcal{F}$}, and so it is a subalgebra of ${\rm End}({\mathcal{F}})$.
Fix $d^{\prime}$$\in{\mathbb{N}}$ and $p^{\prime},q^{\prime},m^{\prime},n^{\prime}\in{\mathbb{Z}}_{+}$ with ${d^{\prime}}\leq d$, $p^{\prime}\leq p$, $q^{\prime}\leq q$, $m^{\prime}\leq m$, and $n^{\prime}\leq n$. Let
${\boldsymbol{\ell}}=p^{\prime}+q^{\prime}+m^{\prime}+n^{\prime}.$
Fix two sequences
${\bf w}:=(w_{1},\ldots,w_{{d^{\prime}}})\in{\mathbb{C}}^{{d^{\prime}}}\qquad\text{and}\qquad{\bf z}:=(z_{1},\ldots,z_{{\boldsymbol{\ell}}})\in{\mathbb{C}}^{{\boldsymbol{\ell}}}.$
We write $d$ (resp., $p$, $q$, $m$ and $n$) into a sum of $d^{\prime}$} (resp., $p^{\prime}$, $q^{\prime}$, $m^{\prime}$ and $n^{\prime}$) positive integers. More precisely, we let $\xi_{1},\ldots,\xi_{{d^{\prime}}}\in{\mathbb{N}}$ be such that
$\sum_{i=1}^{{d^{\prime}}}\xi_{i}=d,$
and let $\gamma_{1},\ldots,\gamma_{{\boldsymbol{\ell}}}\in{\mathbb{N}}$ be such that
(4.1) $\sum_{i=1}^{p^{\prime}}\gamma_{i}=p,\quad\sum_{i=1}^{q^{\prime}}\gamma_{p^{\prime}+i}=q,\quad\sum_{i=1}^{m^{\prime}}\gamma_{p^{\prime}+q^{\prime}+i}=m,\quad\sum_{i=1}^{n^{\prime}}\gamma_{p^{\prime}+q^{\prime}+m^{\prime}+i}=n.$
Set
${\boldsymbol{\xi}}=(\xi_{1},\ldots,\xi_{{d^{\prime}}})\qquad\text{and}\qquad{\boldsymbol{\gamma}}=(\gamma_{1},\ldots,\gamma_{{\boldsymbol{\ell}}}).$
Also, we define $d_{1},\ldots,d_{{d^{\prime}}+1}\in{\mathbb{Z}}_{+}$ by
$d_{1}=0\qquad\text{and}\qquad d_{i+1}=\sum_{j=1}^{i}\xi_{j},\quad 1\leq i\leq{d^{\prime}},$
and define $l_{1},\ldots,l_{{\boldsymbol{\ell}}+1}\in{\mathbb{Z}}_{+}$ by
$l_{1}=0\qquad\text{and}\qquad l_{i+1}=\sum_{j=1}^{i}\gamma_{j},\quad 1\leq i\leq{\boldsymbol{\ell}}.$
For $1\leq i\leq p^{\prime}+q^{\prime}$ and $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$, let
${\bf y}_{(i)}=\{y^{a}_{r}\,|\,l_{i}+1\leq r\leq l_{i+1},1\leq a\leq d\}$
and
${\bf x}_{(j)}=\{x^{a}_{r}\,|\,l_{j}+1-p-q\leq r\leq l_{j+1}-p-q,1\leq a\leq d\}.$
We have
(4.2) ${\mathcal{F}}=\left(\bigotimes_{i=1}^{p^{\prime}+q^{\prime}}{\mathbb{C}}\big[{\bf y}_{(i)}\big]\right)\otimes\left(\bigotimes_{j=p^{\prime}+q^{\prime}+1}^{{\boldsymbol{\ell}}}{\mathbb{C}}\big[{\bf x}_{(j)}\big]\right)\!.$
The following proposition induces an action of $U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ on $\mathcal{F}$}.
Proposition 4.1.
There is a Lie superalgebra homomorphism $\phi:{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\longrightarrow\mathcal{D}$ defined by
(4.3) $\begin{array}[]{cll}e_{ab}\otimes\overline{t}_{z_{i}}^{k}\mapsto-\sum_{r=l_{i}+1}^{l_{i+1}-k}y^{b}_{r}\partial_{y^{a}_{r+k}},\\ e_{ab}\otimes\overline{t}_{z_{j}}^{l}\mapsto\sum_{r=l_{j}+1-p-q}^{l_{j+1}-l-p-q}(-1)^{|p+q+r|}\partial_{x^{b}_{r}}x^{a}_{r+l},\end{array}$}
for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $1\leq i\leq p^{\prime}+q^{\prime}$, $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$ and $1\leq a,b\leq d$. Thus, $\phi$ extends to a superalgebra homomorphism $\phi:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)\longrightarrow\mathcal{D}$.
Proof.
It suffices to verify that $\phi([A,B])=[\phi(A),\phi(B)]$ for all elements $A$ and $B$ on the left-hand side of (4.3). Clearly, for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $1\leq i\leq p^{\prime}+q^{\prime}$, $p^{\prime}+q^{\prime}+1\leq j\leq{\boldsymbol{\ell}}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$,
$\phi\!\left(\left[e_{ab}\otimes\overline{t}_{z_{i}}^{k},e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\!\right]\right)=0=\left[\phi\!\left(e_{ab}\otimes\overline{t}_{z_{i}}^{k}\right),\phi\!\left(e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\right)\!\right].$
Now for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $1\leq i,j\leq p^{\prime}+q^{\prime}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$, we have
$\phi\!\left(\left[e_{ab}\otimes\overline{t}_{z_{i}}^{k},e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\!\right]\right)$ $=$ $\delta_{i,j}\phi\!\left(\left[e_{ab},e_{a^{\prime}b^{\prime}}\!\right]\otimes\overline{t}_{z_{i}}^{k+l}\right)$
$=$ $\delta_{i,j}\phi\!\left((\delta_{a^{\prime},b}e_{ab^{\prime}}-\delta_{a,b^{\prime}}e_{a^{\prime}b})\otimes\overline{t}_{z_{i}}^{k+l}\right)$
$=$ $\delta_{i,j}\left(-\delta_{a^{\prime},b}\sum_{r=l_{i}+1}^{l_{i+1}-k-l}y^{b^{\prime}}_{r}\partial_{y^{a}_{r+k+l}}+\delta_{a,b^{\prime}}\sum_{r=l_{i}+1}^{l_{i+1}-k-l}y^{b}_{r}\partial_{y^{a^{\prime}}_{r+k+l}}\right).$
Meanwhile,
$\left[\phi(e_{ab}\otimes\overline{t}_{z_{i}}^{k}),\phi(e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l})\!\right]$
$\hskip 42.67912pt=\left[\sum_{r=l_{i}+1}^{l_{i+1}-k}y^{b}_{r}\partial_{y^{a}_{r+k}},\sum_{s=l_{j}+1}^{l_{j}+\gamma_{j}-l}y^{b^{\prime}}_{s}\partial_{y^{a^{\prime}}_{s+l}}\!\right]$
$\hskip 42.67912pt=\delta_{i,j}\sum_{r=l_{i}+1}^{l_{i+1}-k}\sum_{s=l_{i}+1}^{l_{i+1}-l}\left[y^{b}_{r}\partial_{y^{a}_{r+k}},y^{b^{\prime}}_{s}\partial_{y^{a^{\prime}}_{s+l}}\,\!\right]$
$\hskip 42.67912pt=\delta_{i,j}\sum_{r=l_{i}+1}^{l_{i+1}-k}\sum_{s=l_{i}+1}^{l_{i+1}-l}\left(y^{b}_{r}\partial_{y^{a}_{r+k}}y^{b^{\prime}}_{s}\partial_{y^{a^{\prime}}_{s+l}}-y^{b^{\prime}}_{s}\partial_{y^{a^{\prime}}_{s+l}}y^{b}_{r}\partial_{y^{a}_{r+k}}\right)$
$\hskip 42.67912pt=\delta_{i,j}\sum_{r=l_{i}+1}^{l_{i+1}-k}\sum_{s=l_{i}+1}^{l_{i+1}-l}\left(y^{b}_{r}\left[\partial_{y^{a}_{r+k}},y^{b^{\prime}}_{s}\,\!\right]\partial_{y^{a^{\prime}}_{s+l}}-y^{b^{\prime}}_{s}\left[\partial_{y^{a^{\prime}}_{s+l}},y^{b}_{r}\,\!\right]\partial_{y^{a}_{r+k}}\right)$
$\hskip 42.67912pt=\delta_{i,j}\left(\sum_{r=l_{i}+1}^{l_{i+1}-k-l}\delta_{a,b^{\prime}}y^{b}_{r}\partial_{y^{a^{\prime}}_{r+k+l}}-\sum_{s=l_{i}+1}^{l_{i+1}-k-l}\delta_{a^{\prime},b}y^{b^{\prime}}_{s}\partial_{y^{a}_{s+k+l}}\right).$
This gives
$\phi\!\left(\left[e_{ab}\otimes\overline{t}_{z_{i}}^{k},e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\!\right]\right)=\left[\phi\!\left(e_{ab}\otimes\overline{t}_{z_{i}}^{k}),\phi(e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\right)\!\right].$
We can show similarly that for $0\leq k\leq\gamma_{i}-1$, $0\leq l\leq\gamma_{j}-1$, $p^{\prime}+q^{\prime}+1\leq i,j\leq{\boldsymbol{\ell}}$ and $1\leq a,b,a^{\prime},b^{\prime}\leq d$,
$\phi\!\left(\left[e_{ab}\otimes t_{z_{i}}^{k},e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\!\right]\right)=\left[\phi\!\left(e_{ab}\otimes t_{z_{i}}^{k}\right),\phi\!\left(e_{a^{\prime}b^{\prime}}\otimes\overline{t}_{z_{j}}^{l}\right)\!\right].$
This completes the proof. ∎
Recall the symbol $\diamondsuit={p+m|q+n}$. For $1\leq a\leq{d^{\prime}}$, define
(4.4) ${\boldsymbol{\Sigma}}^{(a)}=\left\{\,x^{\alpha}_{i},y^{\alpha}_{r}\;\middle|\;d_{a}+1\leq\alpha\leq d_{a+1},1\leq i\leq m+n,1\leq r\leq p+q\,\right\}.$
We have
(4.5) ${\mathcal{F}}=\bigotimes_{a=1}^{d^{\prime}}{\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}].$
The following proposition induces an action of $U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)$ on $\mathcal{F}$}.
Proposition 4.2.
There is a Lie superalgebra homomorphism $\varphi:{\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\longrightarrow\mathcal{D}$ defined by
(4.6) $\begin{array}[]{cll}E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s},\\ E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}},\\ E^{p+q+i}_{s}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}x^{\alpha+k}_{i}y^{\alpha}_{s},\\ E^{p+q+i}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}&\mapsto\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}x^{\alpha+k}_{i}\partial_{x^{\alpha}_{j}},\end{array}$}
for $0\leq k\leq\xi_{a}-1$, $1\leq a\leq{d^{\prime}}$, $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$. Thus, $\varphi$ extends to a superalgebra homomorphism $\varphi:U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)\longrightarrow\mathcal{D}$.
Proof.
We need to verify that $\varphi([A,B])=\left[\,\varphi(A),\varphi(B)\,\!\right]$ for all elements $A$ and $B$ on the left-hand side of (4.6). Clearly, for $0\leq k\leq\xi_{a}-1$, $0\leq l\leq\xi_{b}-1$, $1\leq a,b\leq{d^{\prime}}$, $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$,
$\varphi\!\left(\left[E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k},E^{p+q+i}_{p+q+j}\otimes\overline{t}_{w_{a}}^{l}\!\right]\right)=0=\left[\varphi\!\left(E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{p+q+i}_{p+q+j}\otimes\overline{t}_{w_{a}}^{l}\right)\!\right].$
We will only verify that for $0\leq k\leq\xi_{a}-1$, $0\leq l\leq\xi_{b}-1$, $1\leq a,b\leq{d^{\prime}}$, $1\leq r,s,s^{\prime}\leq p+q$ and $1\leq i,j\leq m+n$,
(4.7) $\varphi\!\left(\left[E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k},E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)=\left[\varphi\!\left(E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\right)\!\right]$
and
(4.8) $\varphi\!\left(\left[E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k},E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)=\left[\varphi\!\left(E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}\right),\varphi\!\left(E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\right)\!\right]\!.$
The verification of all other identities can be carried out analogously. Write $|\overline{i}|=|p+q+i|$ for $1\leq i\leq m+n$. Let us prove (4.7). We have
$\varphi\!\left(\left[E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k},E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)$ $=$ $\delta_{a,b}\varphi\!\left(\left[E^{r}_{s},E^{s^{\prime}}_{p+q+j}\!\right]\otimes\overline{t}_{w_{a}}^{k+l}\right)$
$=$ $\delta_{a,b}\delta_{s,s^{\prime}}\varphi\!\left(E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k+l}\right)$
$=$ $\delta_{a,b}\delta_{s,s^{\prime}}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k-l}\partial_{y^{\alpha+k+l}_{r}}\partial_{x^{\alpha}_{j}}.$
Meanwhile,
$\left[\varphi(E^{r}_{s}\otimes\overline{t}_{w_{a}}^{k}),\varphi(E^{s^{\prime}}_{p+q+j}\otimes\overline{t}_{w_{b}}^{l})\!\right]$
$=$ $\left[\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}(-1)^{|s|+1}\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s},\sum_{\beta=d_{b}+1}^{d_{b+1}-l}\partial_{y^{\beta+l}_{s^{\prime}}}\partial_{x^{\beta}_{j}}\!\right]$
$=$ $\delta_{a,b}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\sum_{\beta=d_{a}+1}^{d_{a+1}-l}(-1)^{|s|+1}\left[\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s},\partial_{y^{\beta+l}_{s^{\prime}}}\partial_{x^{\beta}_{j}}\!\right]$
$=$ $\delta_{a,b}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\sum_{\beta=d_{a}+1}^{d_{a+1}-l}(-1)^{|s|+1}\left(\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s}\partial_{y^{\beta+l}_{s^{\prime}}}\partial_{x^{\beta}_{j}}-(-1)^{(|r|+|s|)(|s^{\prime}|+|\overline{j}|)}\partial_{y^{\beta+l}_{s^{\prime}}}\partial_{x^{\beta}_{j}}\partial_{y^{\alpha+k}_{r}}y^{\alpha}_{s}\right)$
$=$ $\delta_{a,b}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\sum_{\beta=d_{a}+1}^{d_{a+1}-l}(-1)^{|s|+1}\left(\partial_{y^{\alpha+k}_{r}}\left[y^{\alpha}_{s},\partial_{y^{\beta+l}_{s^{\prime}}}\!\right]\partial_{x^{\beta}_{j}}\right)$
$=$ $\delta_{a,b}\delta_{s,s^{\prime}}\sum_{\beta=d_{a}+1}^{d_{a+1}-k-l}\partial_{y^{\beta+k+l}_{r}}\partial_{x^{\beta}_{j}}.$
This completes the proof of (4.7). Now we prove (4.8). We find that
$\varphi\!\left(\left[E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k},E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l}\!\right]\right)$
$=$ $\delta_{a,b}\varphi\!\left(\left[E^{r}_{p+q+j},E^{p+q+i}_{s}\,\!\right]\otimes\overline{t}_{w_{a}}^{k+l}\right)$
$=$ $\delta_{a,b}\varphi\!\left(\left(\delta_{i,j}E^{r}_{s}-(-1)^{(|r|+|\overline{j}|)(|s|+|\overline{i}|)}\delta_{r,s}E^{p+q+i}_{p+q+j}\right)\otimes\overline{t}_{w_{a}}^{k+l}\right)$
$=$ $\delta_{a,b}\Big(\delta_{i,j}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k-l}(-1)^{|s|+1}\partial_{y^{\alpha+k+l}_{r}}y^{\alpha}_{s}-(-1)^{(|r|+|\overline{j}|)(|s|+|\overline{i}|)}\delta_{r,s}\sum_{\beta=d_{a}+1}^{d_{a+1}-k-l}x^{\beta+k+l}_{i}\partial_{x^{\beta}_{j}}\Big).$
Meanwhile,
$\left[\varphi(E^{r}_{p+q+j}\otimes\overline{t}_{w_{a}}^{k}),\varphi(E^{p+q+i}_{s}\otimes\overline{t}_{w_{b}}^{l})\!\right]$
$=$ $\left[\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}},\sum_{\beta=d_{b}+1}^{d_{b+1}-l}(-1)^{|s|+1}x^{\beta+l}_{i}y^{\beta}_{s}\!\right]$
$=$ $\delta_{a,b}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\sum_{\beta=d_{a}+1}^{d_{a+1}-l}(-1)^{|s|+1}\left[\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}},x^{\beta+l}_{i}y^{\beta}_{s}\!\right]$
$=$ $\delta_{a,b}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\sum_{\beta=d_{a}+1}^{d_{a+1}-l}(-1)^{|s|+1}\left(\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}}x^{\beta+l}_{i}y^{\beta}_{s}-(-1)^{(|r|+|\overline{j}|)(|s|+|\overline{i}|)}x^{\beta+l}_{i}y^{\beta}_{s}\partial_{y^{\alpha+k}_{r}}\partial_{x^{\alpha}_{j}}\right)$
$=$ $\delta_{a,b}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k}\sum_{\beta=d_{a}+1}^{d_{a+1}-l}(-1)^{|s|+1}\Big(\partial_{y^{\alpha+k}_{r}}\left[\partial_{x^{\alpha}_{j}},x^{\beta+l}_{i}\,\!\right]y^{\beta}_{s}$
$\hskip 170.71652pt-(-1)^{(|r|+|\overline{j}|)(|s|+|\overline{i}|)}x^{\beta+l}_{i}\left[y^{\beta}_{s},\partial_{y^{\alpha+k}_{r}}\!\right]\partial_{x^{\alpha}_{j}}\Big)$
$=$ $\!\delta_{a,b}\Big(\delta_{i,j}\sum_{\beta=d_{a}+1}^{d_{a+1}-k-l}(-1)^{|s|+1}\partial_{y^{\beta+k+l}_{r}}y^{\beta}_{s}$
$\hskip 128.0374pt-(-1)^{(|r|+|\overline{j}|)(|s|+|\overline{i}|)}\delta_{r,s}\sum_{\alpha=d_{a}+1}^{d_{a+1}-k-l}x^{\alpha+k+l}_{i}\partial_{x^{\alpha}_{j}}\Big)\!.$
This completes the proof of (4.8). ∎
The pair $(\phi,\varphi)$ of homomorphisms induce a joint action of $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on the Fock space $\mathcal{F}$}. We will see, in Section 4.3, that $(\phi,\varphi)$ gives a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$, which is an equivalence between the actions of $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{{p+m|q+n}}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on $\mathcal{F}$}.
Remark 4.3.
In the special case where ${\boldsymbol{\xi}}=(1^{d^{\prime}})$ and ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, a variant of the pair $(\phi,\varphi)$ was considered in [CLZ, pp. 789–790] and [ChL2, Section 4.3]. It gives rise to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$ (see [CLZ, Theorem 3.3]).
4.2. Examples
In this subsection, we will give examples of ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$-modules that will be used in Section 4.4.
Let
${\bf d}_{{\mathcal{F}}}=\sum_{a=1}^{d}\left(-\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}x^{a}_{i}\partial_{x^{a}_{i}}\right)\!.$
For any monomial $f\in{\mathcal{F}}$, we have ${\bf d}_{{\mathcal{F}}}(f)=kf$ for some $k\in{\mathbb{Z}}$, and we define the degree of $f$ to be $k$. In particular, the degree of each $x^{a}_{i}$} is $1$, and the degree of each $y^{a}_{r}$ is 1-1. For this reason, it is natural to call $\bf d$$_{{\mathcal{F}}}$ the degree operator on $\mathcal{F}$}.
By definition, ${\mathfrak{gl}}_{\diamondsuit}^{(d)}[t]={\mathfrak{gl}}_{\diamondsuit}(0,d)$, and by Proposition 4.2, the Fock space $\mathcal{F}$} is a ${\mathfrak{gl}}^{(d)}_{\diamondsuit}[t]$-module. For each $k\in{\mathbb{Z}}$, let $V^{(d)}_{k}$ denote the subspace of $\mathcal{F}$} spanned by monomials of degree $k$. Since $\bf d$$_{{\mathcal{F}}}$ commutes with the action of ${\mathfrak{gl}}^{(d)}_{\diamondsuit}[t]$ on $\mathcal{F}$}, the space $V^{(d)}_{k}$ is a ${\mathfrak{gl}}^{(d)}_{\diamondsuit}[t]$-module. We have a decomposition of $\mathcal{F}$} into a direct sum of ${\mathfrak{gl}}^{(d)}_{\diamondsuit}[t]$-modules
(4.9) ${\mathcal{F}}=\bigoplus_{k\in{\mathbb{Z}}}V^{(d)}_{k}$.
For $1\leq a\leq{d^{\prime}}$, recall ${\boldsymbol{\Sigma}}^{(a)}$ defined in (4.4). The Fock space ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}]$ is a ${\mathfrak{gl}}_{\diamondsuit}({w_{a},\xi_{a}})$-module. Analogous to (4.9), there is also a direct sum decomposition of ${\mathfrak{gl}}_{\diamondsuit}({w_{a},\xi_{a}})$-modules
${\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}]=\bigoplus_{k\in{\mathbb{Z}}}V_{k}^{(\xi_{a})},$
where $V_{k}^{(\xi_{a})}$ is the subspace of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}]$ spanned by monomials of degree $k$. Define
(4.10) $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}=V^{(\xi_{1})}_{k_{1}}\otimes\cdots\otimes V^{(\xi_{{d^{\prime}}})}_{k_{{d^{\prime}}}}\qquad\mbox{for \, $k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}$}.$
By Proposition 4.2 and (4.5), $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}$ is a ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$-module, and we have a direct sum decomposition of ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$-modules
(4.11) ${\mathcal{F}}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}.$
4.3. A duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
The $k\times k$ matrix
${\rm J}_{k}(\lambda):=\begin{bmatrix}\lambda&1&\ldots&0\\ 0&\lambda&\ddots&\vdots\\ \vdots&\ddots&\ddots&1\\ 0&\ldots&0&\lambda\end{bmatrix}$
is called a Jordan block with eigenvalue $\lambda$. If $\lambda\not=0$, then
(4.12) $\big(-{\rm J}_{k}(-\lambda)\big)^{-1}=\begin{bmatrix}\lambda^{-1}&\lambda^{-2}&\ldots&\lambda^{-k}\\ 0&\lambda^{-1}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\lambda^{-2}\\ 0&\ldots&0&\lambda^{-1}\end{bmatrix}$.
Consider the Jordan matrix
${\rm J}_{\boldsymbol{\xi}}({\bf w}):=\bigoplus_{a=1}^{{d^{\prime}}}{\rm J}_{\xi_{a}}(w_{a}):=\begin{bmatrix}{\rm J}_{\xi_{1}}(w_{1})&0&\ldots&0\\ 0&{\rm J}_{\xi_{2}}(w_{2})&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\ldots&0&{\rm J}_{\xi_{d^{\prime}}}(w_{d^{\prime}})\end{bmatrix}$,
which is the direct sum of Jordan blocks. Define $\mu_{{\boldsymbol{\xi}}}^{{\bf w}}\in{\mathfrak{gl}}_{d}^{*}$ by
(4.13) $\mu_{{\boldsymbol{\xi}}}^{{\bf w}}(e_{ab})=-{\rm J}_{\boldsymbol{\xi}}({\bf w})_{a,b}\qquad\mbox{for $a,b=1,\ldots,d$}.$
Here ${\rm J}_{\boldsymbol{\xi}}({\bf w})_{a,b}$ denotes the $(a,b)$ entry of ${\rm J}_{\boldsymbol{\xi}}({\bf w})_{a,b}$.
Let
${\rm J}_{{\boldsymbol{\gamma}}}({\bf z})=\bigoplus_{i=1}^{{\boldsymbol{\ell}}}{\rm J}_{\gamma_{i}}(z_{i})$
Define $\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}\in{\mathfrak{gl}}_{\diamondsuit}^{*}$ by
(4.14) $\hskip 28.45274pt\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}(E^{i}_{j})=(-1)^{|i|+1}{\rm J}_{{\boldsymbol{\gamma}}}({\bf z})_{i,j}\qquad\mbox{for $i,j\in\mathbb{I}$}.$
Note that $\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}$ vanishes on the odd part of $\mathfrak{gl}$$_{\diamondsuit}$.
We write
(4.15) $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\mathcal{A}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\mathcal{A}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$
We also write
$\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\widehat{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad{\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})={\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}),$
where the matrices $\widehat{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})$ and ${\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}})$ are defined as in (3.3) and (3.4).
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{\diamondsuit}$. As discussed earlier, the Gaudin algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ are determined by ${\rm cdet}\,\!\left(\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)$ and ${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)$, respectively.
The maps $\phi$ and $\varphi$, given in Proposition 4.1 and Proposition 4.2, extend naturally to the superalgebra homomorphisms
$\phi:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
and
$\varphi:U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
respectively. For $1\leq i\leq{\boldsymbol{\ell}}$, define
(4.16) $\gamma_{i}]=\begin{cases}\,\gamma_{i},&\mbox{if}\ \ 1\leq i\leq p^{\prime}\quad\text{or}\quad p^{\prime}+q^{\prime}+1\leq i\leq p^{\prime}+q^{\prime}+m^{\prime};\\ -\gamma_{i},&\mbox{otherwise}.\end{cases}$
Theorem 4.4.
We have
$\prod_{i=1}^{{\boldsymbol{\ell}}}(\partial_{z}-z_{i})^{[\gamma_{i}]}\cdot\phi\!\left({\rm cdet}\,\!\left(\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)\right)=\prod_{a=1}^{{d^{\prime}}}(z-w_{a})^{\xi_{a}}\cdot\varphi\!\left({\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)\right)$
as an equality between elements of $\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$.
Proof.
Let
$J=\bigoplus_{a=1}^{{d^{\prime}}}\big(\!-{\rm J}_{\xi_{a}}(w_{a}-z)\big)\quad\text{and}\quad J^{\prime}=\bigoplus_{i=1}^{{\boldsymbol{\ell}}}\big(\!-{\rm J}_{\gamma_{i}}(z_{i}-\partial_{z})\big).$
By (3.2), we find that
(4.17) $\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=J^{t}-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{e_{ab}\otimes\overline{t}_{z_{i}}^{k}}{(\partial_{z}-z_{i})^{k+1}}\!\right]^{t}_{\!a,b=1,\ldots,d}$
and
(4.18) ${\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=J^{\prime}-\left[(-1)^{|i|}\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}}{(z-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}.$
Consider the $(m+n)\times d$ matrices
$X:=\big[(-1)^{|p+q+i|}x^{a}_{i}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}\quad\text{and}\quad P_{X}:=\big[\partial_{x^{a}_{i}}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}$
and the $(p+q)\times d$ matrices
$Y:=\big[(-1)^{|r|+1}y^{a}_{r}\big]^{a=1,\ldots,d}_{r=1,\ldots,p+q}\qquad\text{and}\qquad P_{Y}:=\big[(-1)^{|r|}\partial_{y^{a}_{r}}\big]^{a=1,\ldots,d}_{r=1,\ldots,p+q}.$
Write
$M=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\quad\text{and}\quad M^{\prime}=\begin{bmatrix}P_{Y}\vskip 8.53581pt\\ X\end{bmatrix}$.
Let
$\mathfrak{L}=\begin{bmatrix}J^{t}&M\vskip 5.69054pt\\ M^{\prime}&J^{\prime}\end{bmatrix}$.
It is straightforward to check that $\mathfrak{L}$ is an amply invertible Manin matrix of type $\hat{{\bf s}}:=(0^{d},{\bf s})\in\mathcal{S}_{d+p+m|q+n}$. By Proposition 2.1 and Proposition 2.5, we have
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})={\rm cdet}\,(J^{t})\cdot{\rm Ber}^{\bf s}\,(J^{\prime}-M^{\prime}(J^{t})^{-1}M).$
We see that
$\varphi\!\left({\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)\right)$ $=$ ${\rm Ber}^{\bf s}\,\left(\!J^{\prime}-\left[(-1)^{|i|}\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{\varphi\big(E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}\big)}{(z-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}\right)$
$=$ ${\rm Ber}^{\bf s}\,(J^{\prime}-M^{\prime}(J^{t})^{-1}M).$
Here the first equality follows from (4.18) while the second follows from Proposition 4.2 and (4.12). Consequently,
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=\prod_{a=1}^{{d^{\prime}}}(z-w_{a})^{\xi_{a}}\cdot\varphi\left({\rm Ber}^{\bf s}\,\!\big({\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big)\right)\!.$
On the other hand, as $J^{t}-M(J^{\prime})^{-1}M^{\prime}$ is a sufficiently invertible Manin matrix of type $(0^{d})$, Proposition 2.4 and Proposition 2.5 imply that
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})={\rm Ber}^{\bf s}\,(J^{\prime})\cdot{\rm cdet}\,(J^{t}-M(J^{\prime})^{-1}M^{\prime}).$
By (4.12), (4.17) and Proposition 4.1,
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})$ $=$ $\prod_{i=1}^{{\boldsymbol{\ell}}}(\partial_{z}-z_{i})^{[\gamma_{i}]}\cdot{\rm cdet}\,\left(J^{t}-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{\phi(e_{a,b}\otimes\overline{t}_{z_{i}}^{k})}{(\partial_{z}-z_{i})^{k+1}}\!\right]^{t}_{\!a,b=1,\ldots,d}\right)$
$=$ $\prod_{i=1}^{{\boldsymbol{\ell}}}(\partial_{z}-z_{i})^{[\gamma_{i}]}\cdot\phi\!\left({\rm cdet}\,\!\left(\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)\right).$
This proves the theorem. ∎
We have the following duality between the Gaudin algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$. We call it the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$ for Gaudin models with irregular singularities.
Theorem 4.5.
$\phi\big(\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\varphi\big(\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big).$
Proof.
This follows from Proposition 3.3 and Theorem 4.4. ∎
The Weyl superalgebra $\mathcal{D}$ is a subalgebra of ${\rm End}({\mathcal{F}})$. The maps $\phi$ and $\varphi$ induce actions of $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on $\mathcal{F}$}. Theorem 4.5 also says that these actions are equivalent.
If any of $p,q,m,n$ is set to 0, Theorem 4.5 will give a special version of the duality. Taking $p=q=n=0$, Theorem 4.4 gives the identity in [VY, Theorem 4.8], and we recover the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ due to Vicedo and Young.
Corollary 4.6 ([VY, Theorem 4.8]).
Set $p=q=n=0$. Then $\phi\big(\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\varphi\left(\mathcal{A}_{m}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)$.
Theorem 4.5 also gives a fermionic realization of the above duality.
Corollary 4.7.
Set $p=q=m=0$. Then $\phi\big(\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\varphi\left(\mathcal{A}_{0|n}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)$.
Taking $q=n=0$ in Theorem 4.5, we obtain the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m+p})$ for the bosonic oscillators in which the Fock space $\mathcal{F}$} decomposes into a direct sum of tensor products of infinite-dimensional ${\mathfrak{gl}}_{m+p}$-modules.
Corollary 4.8.
Set $q=n=0$. Then $\phi\big(\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\varphi\left(\mathcal{A}_{m+p}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)$.
Further, the following duality holds.
Corollary 4.9.
Suppose ${d^{\prime}}=d$ and ${\boldsymbol{\ell}}=p+q+m+n$. Then $\phi\!\left(\mathcal{A}_{d}^{{\bf w}\!,\,(1^{d^{\prime}})}\left({\bf z},(1^{\boldsymbol{\ell}})\right)\right)=\varphi\!\left(\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{\boldsymbol{\ell}})}\left({\bf w},(1^{d^{\prime}})\right)\right)$.
Remark 4.10.
A variant of Corollary 4.9, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$, is established in [ChL2].
4.4. An application of the duality
In this subsection, we consider the actions of $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on the Fock space $\mathcal{F}$} via the maps $\phi$ and $\varphi$, respectively; see Proposition 4.1 and Proposition 4.2. We will restrict our attention to ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, where ${\boldsymbol{\ell}}=p+q+m+n$.
First of all, we review some well-known properties of the action of the Gaudin algebras (with the simplest possible singularities) on finite-dimensional ${\mathfrak{gl}}_{d}$-modules. An element $\mu\in{\mathfrak{gl}}_{d}^{*}$ is called regular if the dimension of the centralizer ${\mathfrak{gl}}_{d}^{\mu}$ of $\mu$ in ${\mathfrak{gl}}_{d}$ is $d$. For $\ell\in{\mathbb{N}}$, let
${\boldsymbol{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\}$
be the configuration space of $\ell$ distinct points on $\mathbb{C}$$^{\ell}$. The following is a consequence of [FFRy].
Proposition 4.11.
Let $\underline{V\mkern-4.0mu}\mkern 4.0mu=V_{1}\otimes\cdots\otimes V_{\boldsymbol{\ell}}$, where $V_{1},\ldots,V_{\boldsymbol{\ell}}$ are finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules. Then $\underline{V\mkern-4.0mu}\mkern 4.0mu$ is a cyclic $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))$-module for ${\bf w}\in{\boldsymbol{X}}_{\!{d^{\prime}}}$ and ${\bf z}\in{\boldsymbol{X}}_{\!{\boldsymbol{\ell}}}$. Moreover, the Gaudin algebra $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))_{\underline{V\mkern-4.0mu}\mkern 4.0mu}$ is diagonalizable with a simple spectrum for generic ${\bf w}\in{\boldsymbol{X}}_{\!{d^{\prime}}}$ and ${\bf z}\in{\boldsymbol{X}}_{\!{\boldsymbol{\ell}}}$.
Proof.
The first statement follows from [FFRy, Corollary 5] since the element $\mu_{{\boldsymbol{\xi}}}^{{\bf w}}\in{\mathfrak{gl}}_{d}^{*}$, defined in (4.13), is regular. The second statement follows from the first one, [FFRy, Lemma 2], and the property that having a simple spectrum is an open condition. ∎
[…]
(4.19) $M_{\mu}=\left\{\,\!v\in M\;\middle|\;hv=\mu(h)v\quad\mbox{for all $h\in{\mathfrak{h}}_{\diamondsuit}$}\!\,\right\}.$
If $M_{\mu}\not=0$, then $\mu$ is called a weight of $M$ and $M_{\mu}$ is called the $\mu$-weight space of $M$. There is a weight space decomposition
(4.20) ${\mathcal{F}}=\bigoplus_{\mu\in{\mathfrak{h}}_{\diamondsuit}^{*}}{\mathcal{F}}_{\mu}.$
Let ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, where ${\boldsymbol{\ell}}=p+q+m+n$. For $1\leq i\leq{\boldsymbol{\ell}}$, define ${\mathfrak{gl}}_{d}^{(i)}={\mathfrak{gl}}_{d}(z_{i},1)\cong{\mathfrak{gl}}_{d}$. Let $W^{(i)}_{\mu_{i}}$ be the subspace of ${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]$ (resp., ${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]$) spanned by monomials of degree $\mu_{i}\in-{\mathbb{Z}}_{+}$ (resp., $\mathbb{Z}$$_{+}$) if $1\leq i\leq p+q$ (resp., $p+q+1\leq i\leq{\boldsymbol{\ell}}$). We have direct sum decompositions of ${\mathfrak{gl}}_{d}^{(i)}$-modules:
${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]=\bigoplus_{\mu_{i}\in-{\mathbb{Z}}_{+}}W^{(i)}_{\mu_{i}},\qquad 1\leq i\leq p+q,$
and
${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]=\bigoplus_{\mu_{i}\in{\mathbb{Z}}_{+}}W^{(i)}_{\mu_{i}},\qquad p+q+1\leq i\leq{\boldsymbol{\ell}}.$
Evidently, $W^{(i)}_{\mu_{i}}$ is finite-dimensional, and it is well known that $W^{(i)}_{\mu_{i}}$ is an irreducible ${\mathfrak{gl}}_{d}^{(i)}$-module for each $i$ (see, for instance, [CLZ, Theorem 3.3] by setting exactly one of $p,q,m,n$ (given there) to 1 and the others to 0).
For $\mu_{1},\ldots,\mu_{p+q}\in-{\mathbb{Z}}_{+}$ and $\mu_{p+q+1},\ldots,\mu_{\boldsymbol{\ell}}\in{\mathbb{Z}}_{+}$, define
$\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}=W_{\mu_{1}}^{(1)}\otimes\cdots\otimes W_{\mu_{\boldsymbol{\ell}}}^{({\boldsymbol{\ell}})}.$
The Fock space $\mathcal{F}$} decomposes into a direct sum of ${\mathfrak{gl}}_{d}({\bf z},(1^{\boldsymbol{\ell}}))$-modules
(4.21) ${\mathcal{F}}=\bigoplus_{\begin{subarray}{c}\mu_{1},\ldots,\mu_{p+q}\in-{\mathbb{Z}}_{+},\\ \mu_{p+q+1},\ldots,\mu_{\boldsymbol{\ell}}\in{\mathbb{Z}}_{+}\end{subarray}}\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})},$
and each $\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}$ is a ${\mathfrak{gl}}_{d}$-module via the diagonal map from ${\mathfrak{gl}}_{d}$ to ${\mathfrak{gl}}_{d}({\bf z},(1^{\boldsymbol{\ell}}))$.
For any ${\mathfrak{gl}}_{d}$-module $M$ and $\mu\in{\mathfrak{h}}_{d}^{*}$, the space $M_{\mu}$ is defined as in (4.19) by replacing $\mathfrak{h}$$_{\diamondsuit}$ with ${\mathfrak{h}}_{d}$. The action of ${\mathfrak{h}}_{d}$ on $\mathcal{F}$} is determined by the action of $e_{aa}$ on $\mathcal{F}$}, which is given by
$\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}(-1)^{|p+q+i|}\partial_{x^{a}_{i}}x^{a}_{i}=-\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}x^{a}_{i}\partial_{x^{a}_{i}}+(m-n),$
for $1\leq a\leq d$. For $k_{1},\ldots k_{{d^{\prime}}}\in{\mathbb{Z}}$, we define
${\mathcal{F}}_{[k_{1},\ldots,k_{d^{\prime}}]}=\bigoplus_{\eta}{\mathcal{F}}_{\eta},$
where the direct sum is taken over all $\eta\in{\mathfrak{h}}_{d}^{*}$ such that $\sum_{i=d_{a}+1}^{d_{a+1}}\eta(e_{ii})=k_{a}+(m-n)\xi_{a}$ for $1\leq a\leq{d^{\prime}}$. Thus, $\mathcal{F}$} has a direct sum decomposition of subspaces
(4.22) ${\mathcal{F}}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}{\mathcal{F}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$
For $\mu_{1},\ldots,\mu_{p+q}\in-{\mathbb{Z}}_{+}$ and $\mu_{p+q+1},\ldots,\mu_{\boldsymbol{\ell}}\in{\mathbb{Z}}_{+}$, let
$\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}_{[k_{1},\ldots,k_{{d^{\prime}}}]}=\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}\cap{\mathcal{F}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$
Then
(4.23) $\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$
For $k_{1},\ldots k_{{d^{\prime}}}\in{\mathbb{Z}}$, recall that $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}=V^{(\xi_{1})}_{k_{1}}\otimes\cdots\otimes V^{(\xi_{{d^{\prime}}})}_{k_{{d^{\prime}}}}$, where $V^{(\xi_{a})}_{k_{a}}$ is the ${\mathfrak{gl}}_{\diamondsuit}(w_{a},\xi_{a})$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{(a)}]$ spanned by monomials of degree $k_{a}$ (see Section 4.2). Note that $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}$ is infinite-dimensional if $p\not=0$ and $m\not=0$.
Proposition 4.12.
For any $k_{1},\ldots k_{{d^{\prime}}}\in{\mathbb{Z}}$ and any weight $\mu$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}$, we have
$\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}={\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]},$
where
$\overline{\mu}_{r}=\begin{cases}\,\mu(E^{r}_{r})+(-1)^{|r|}d&\mbox{if}\ \ 1\leq r\leq p+q;\\ \,\mu(E^{r}_{r})\ \ &\mbox{if}\ \ p+q+1\leq r\leq{\boldsymbol{\ell}}.\end{cases}$
Proof.
Let $k_{1},\ldots k_{{d^{\prime}}}\in{\mathbb{Z}}$. By (4.11) and (4.22), we have
${\mathcal{F}}_{[k_{1},\ldots,k_{d^{\prime}}]}=\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}.$
The Cartan subalgebra $\mathfrak{h}$$_{\diamondsuit}$ of $\mathfrak{gl}$$_{\diamondsuit}$ acts on $\mathcal{F}$} as follows: For $1\leq r\leq p+q$, $E^{r}_{r}$ acts on $\mathcal{F}$} by
$\sum_{a=1}^{{d^{\prime}}}\sum_{\alpha=d_{a}+1}^{d_{a+1}}(-1)^{|r|+1}\partial_{y^{\alpha}_{r}}y^{\alpha}_{r}=-\sum_{\alpha=1}^{d}y^{\alpha}_{r}\partial_{y^{\alpha}_{r}}+(-1)^{|r|+1}d,$
and for $1\leq i\leq m+n$, $E^{p+q+i}_{p+q+i}$ acts on $\mathcal{F}$} by
$\sum_{a=1}^{{d^{\prime}}}\sum_{\alpha=d_{a}+1}^{d_{a+1}}x^{\alpha}_{i}\partial_{x^{\alpha}_{i}}=\sum_{\alpha=1}^{d}x^{\alpha}_{i}\partial_{x^{\alpha}_{i}}.$
For any weight $\mu$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}$, the above descriptions together with (4.20) and (4.21) give
${\mathcal{F}}_{\mu}={\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}.$
Hence
$\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}={\mathcal{F}}_{[k_{1},\ldots,k_{d^{\prime}}]}\cap{\mathcal{F}}_{\mu}={\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]},$
as desired. ∎
According to the duality for Gaudin models with irregular singularities, we expect that an analog of Proposition 4.11 should be valid for the action of the Gaudin algebra $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$ on some ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$-modules. The following theorem supports our expectation.
Theorem 4.13.
For any $k_{1},\ldots k_{{d^{\prime}}}\in{\mathbb{Z}}$ and any weight $\mu$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}$, the $\mu$-weight space $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}$ is a cyclic $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$-module for ${\bf w}\in{\boldsymbol{X}}_{\!{d^{\prime}}}$ and ${\bf z}\in{\boldsymbol{X}}_{\!{\boldsymbol{\ell}}}$. Moreover, the Gaudin algebra $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})_{\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\!\mu}}$ is diagonalizable with a simple spectrum for generic $\bf w$} and $\bf z$}.
Proof.
Let $k_{1},\ldots k_{{d^{\prime}}}\in{\mathbb{Z}}$, and let $\mu$ be any weight of $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}$. The Gaudin algebra $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$ commutes with the diagonal action of the centralizer ${\mathfrak{gl}}_{\diamondsuit}^{\nu}$ of $\nu$ in $\mathfrak{gl}$$_{\diamondsuit}$ (see [Ry, Proposition 4]), where $\nu:=\nu_{(1^{\boldsymbol{\ell}})}^{{\bf z}}$ is defined as in (4.14). Since $\nu$ corresponds to a diagonal matrix, ${\mathfrak{gl}}_{\diamondsuit}^{\nu}$ contains $\mathfrak{h}$$_{\diamondsuit}$. Thus, $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$ preserves weight spaces, and the weight space $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}$ is an $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})$-module. By Theorem 4.5 and Proposition 4.12, ${\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}$ is an $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},1^{\ell})$-module, and
(4.24) $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})_{\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}}=\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))_{{\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}}.$
As in (4.23), $\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}$} is a direct sum decomposition of $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))$-modules. We apply Proposition 4.11 to the decomposition of $\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}$} to see that the space ${\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}$ is a cyclic $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))$-module for ${\bf w}\in{\boldsymbol{X}}_{\!{d^{\prime}}}$ and ${\bf z}\in{\boldsymbol{X}}_{\!{\boldsymbol{\ell}}}$, and $\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))_{{\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}}$ is diagonalizable with a simple spectrum for generic $\bf w$} and $\bf z$}. In view of (4.24), the theorem follows. ∎
5. A duality for classical Gaudin models
This section is devoted to introducing the classical Gaudin algebras with irregular singularities for ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ and establishing a duality for classical Gaudin models.
5.1. Settings
For any Lie superalgebra $\mathfrak{g}$}, the universal enveloping algebra $U({\mathfrak{g}})$ of $\mathfrak{g}$} has a filtration
(5.1) ${\mathbb{C}}=U_{0}({\mathfrak{g}})\subset U_{1}({\mathfrak{g}})\subset U_{2}({\mathfrak{g}})\subset\cdots U_{i}({\mathfrak{g}})\subset\cdots$
such that $\bigcup_{i=0}^{\infty}U_{i}({\mathfrak{g}})=U({\mathfrak{g}})$ and $U_{i}({\mathfrak{g}})U_{j}({\mathfrak{g}})\subseteq U_{i+j}({\mathfrak{g}})$ for all $i,j\in{\mathbb{Z}}_{+}$, where $U_{i}({\mathfrak{g}})$ is the subspace of $U({\mathfrak{g}})$ spanned by all products $A_{1}\ldots A_{j}$, for $j\leq i$ and $A_{k}\in{\mathfrak{g}}$. We call (5.1) the canonical filtration on $U({\mathfrak{g}})$.
Let ${\rm gr}(U({\mathfrak{g}}))=\bigoplus_{i=0}^{\infty}U_{i}({\mathfrak{g}})/U_{i-1}({\mathfrak{g}})$ (where $U_{-1}({\mathfrak{g}}):=0$). It is called the associated graded algebra of $U({\mathfrak{g}})$ with respect to the canonical filtration. Let $\mathcal{S}({\mathfrak{g}})$ denote the supersymmetric algebra of $\mathfrak{g}$}. We have the following superalgebra isomorphism (see [Dix, Proposition 2.3.6]):
(5.2) ${\rm gr}\!\left(U({\mathfrak{g}})\right)\cong\mathcal{S}({\mathfrak{g}}).$
Because of (5.2), we say that $U({\mathfrak{g}})$ is a quantization of $\mathcal{S}({\mathfrak{g}})$.
A Poisson superalgebra $\mathcal{A}$ is a supercommutative superalgebra over $\mathbb{C}$} equipped with a $\mathbb{C}$}-bilinear map $\{\cdot,\cdot\}:\mathcal{A}\times\mathcal{A}\longrightarrow\mathcal{A}$, called a Poisson bracket, such that $(\mathcal{A},\{\cdot,\cdot\})$ is a Lie superalgebra satisfying the Leibniz rule, i.e.,
$\{A,BC\}=\{A,B\}C+(-1)^{|A||B|}B\{A,C\}$
for any homogeneous elements $A,B,C\in\mathcal{A}$.
The superalgebra ${\rm gr}\!\left(U({\mathfrak{g}})\right)$ is naturally a Poisson superalgebra (the construction in the proof of [CG, Proposition 1.3.2] works here by taking into account the sign rule). This induces a Poisson superalgebra structure on $\mathcal{S}({\mathfrak{g}})$ via (5.2).
Let $\mathfrak{g}$} be a general linear Lie (super)algebra. Fix $\ell\in{\mathbb{N}}$, $\bf z$$\in{\mathbb{C}}^{\ell}$ and $\boldsymbol{\gamma}$$\in{\mathbb{N}}^{\ell}$. For even variables $t$ and $z$, the map $\Psi_{({\bf z},{\boldsymbol{\gamma}})}:=\Psi^{0}_{({\bf z},{\boldsymbol{\gamma}})}$, defined in Section 3.3, preserves filtrations and hence descends to the Poisson superalgebra homomorphism
$\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}:\mathcal{S}(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow\mathcal{S}\big({\mathfrak{g}}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt].$
We add $z$ to the symbol $\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}$ to emphasize the dependency on $z$. Also,
(5.3) $\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}\left(A\otimes t^{-1}\right)=-\sum_{i=1}^{\ell}\sum_{k=0}^{\gamma_{i}-1}\frac{A\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}},\qquad\mbox{for $A\in{\mathfrak{g}}$.}$
We will introduce the classical Gaudin algebras for ${\mathfrak{gl}}_{d}$ and $\mathfrak{gl}$$_{\diamondsuit}$ separately. From now on, we fix two commuting even variables $z$ and $w$.
5.2. Classical Gaudin algebras for ${\mathfrak{gl}}_{d}$
Recall the definition of the determinant given in Section 2.2. Let $\bf z$$\in{\mathbb{C}}^{\ell}$ and $\boldsymbol{\gamma}$$\in{\mathbb{N}}^{\ell}$. For any $\mu\in{\mathfrak{gl}}_{d}^{*}$, we consider the matrix
$\overline{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\delta_{i,j}w+\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}\left(e_{ij}\otimes t^{-1}\right)+\mu(e_{ij})\Big]_{i,j=1,\ldots,d}$
over the commutative algebra $\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt][w]$. We have an expansion
$\det(\overline{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}}))=w^{d}+\sum_{i=1}^{d}\overline{a}_{i}(z)w^{i-1}$
for some $\overline{a}_{i}(z)\in\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt]$. Let $\overline{\mathcal{A}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of the series $\overline{a}_{i}(z)$ for $i=1,\ldots,d$. It is called the classical Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$, or simply the classical Gaudin algebra corresponding to $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$. The algebra $\overline{\mathcal{A}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ is known to be Poisson-commutative [FFTL].
Let ${\rm gr}\!\left(\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)$ be the associated graded algebra of $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ with respect the filtration induced by the canonical filtration on $U({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}}))$. We remark that the proof of [FFTL, Theorem 3.4] shows that $\mathcal{A}_{d}^{0}({\bf z},{\boldsymbol{\gamma}})$ is a quantization of $\overline{\mathcal{A}}_{d}^{0}({\bf z},{\boldsymbol{\gamma}})$, i.e.,
(5.4) ${\rm gr}\!\left(\mathcal{A}_{d}^{0}({\bf z},{\boldsymbol{\gamma}})\right)\cong\overline{\mathcal{A}}_{d}^{0}({\bf z},{\boldsymbol{\gamma}}).$
For ${\bf z}=(0)$ and ${\boldsymbol{\gamma}}=(1)$, the algebra $\overline{\mathcal{A}}_{d}^{\mu}:=\overline{\mathcal{A}}_{d}^{\mu}((0),(1))$ is simply the shift of argument subalgebra of $\mathcal{S}({\mathfrak{g}})$ introduced by Mishchenko and Fomenko; see [MF]. Let $\mathcal{A}_{d}^{\mu}=\mathcal{A}_{d}^{\mu}((0),(1))$. Futorny and Molev [FM, Main Theorem] show that
(5.5) $\hskip 28.45274pt{\rm gr}(\mathcal{A}_{d}^{\mu})\cong\overline{\mathcal{A}}_{d}^{\mu}\quad\qquad\mbox{for any $\mu\in{\mathfrak{gl}}_{d}^{*}$.}$
We expect a more general statement to hold; see Conjecture 5.1 below.
5.3. Classical Gaudin algebras for ${\mathfrak{gl}}_{p+m|q+n}$
Recall the symbol $\diamondsuit={p+m|q+n}$. Let $\bf z$$\in{\mathbb{C}}^{\ell}$ and $\boldsymbol{\gamma}$$\in{\mathbb{N}}^{\ell}$. For any $\mu\in{\mathfrak{gl}}_{\diamondsuit}^{*}$ which vanishes on the odd part of $\mathfrak{gl}$$_{\diamondsuit}$, let
$\overline{{\mathcal{L}}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\delta_{i,j}z+(-1)^{|i|}\left(\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,w}\left(E^{i}_{j}\otimes t^{-1}\right)+\mu(E^{i}_{j})\right)\Big]_{i,j\in\mathbb{I}},$
which is an amply invertible $(p+q+m+n)\times(p+q+m+n)$ matrix of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$ over the supercommutative superalgebra $\mathcal{S}\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},w^{-1})\!)$. We have an expansion
${\rm Ber}^{\bf s}\,\!\left(\overline{{\mathcal{L}}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)=\sum_{i=-\infty}^{p+m-q-n}\overline{b}_{i}(w)z^{i}$
for some $\overline{b}_{i}(w)\in\mathcal{S}\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(w^{-1})\!)$. Let $\overline{\mathcal{A}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $S\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of the series $\overline{b}_{i}(w)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the classical Gaudin algebra for $\mathfrak{gl}$$_{\diamondsuit}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$, or simply the classical Gaudin algebra corresponding to $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$. It is Poisson-commutative by an argument parallel to the proof of [MR, Corollary 3.6]. If $q=n=0$ and $d=p+m$, the algebra coincides with the one given in Section 5.2.
Let ${\rm gr}\!\left(\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)$ be the associated graded algebra of $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ with respect to the filtration induced by the canonical filtration on $U({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}}))$. We conjecture that the Gaudin algebras are quantizations of the classical Gaudin algebras.
Conjecture 5.1.
${\rm gr}\left(\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)\cong\overline{\mathcal{A}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ for any $\mu\in{\mathfrak{gl}}_{\diamondsuit}^{*}$ which vanishes on the odd part of $\mathfrak{gl}$$_{\diamondsuit}$.
The conjectures hold for some special cases (see (5.4) and (5.5)), but the general case is wide open.
5.4. A classical duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
[…]
Recall the Weyl superalgebra $\mathcal{D}$ defined in Section 4.1. We define the degree of any monomial in $\mathcal{D}$ in the usual way. In particular, the degrees of the generators $x^{a}_{i}$, $y^{a}_{r}$, $\partial_{x^{a}_{i}}$ and $\partial_{y^{a}_{r}}$ of $\mathcal{D}$ are defined to be $1$, for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$. For $k\in{\mathbb{N}}$, let $\mathcal{D}_{k}$ be the subspace of $\mathcal{D}$ spanned by all monomials of degree less than or equal to $k$ in the generators $x^{a}_{i}$, $y^{a}_{r}$, $\partial_{x^{a}_{i}}$ and $\partial_{y^{a}_{r}}$, $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$, and let $\mathcal{D}_{0}={\mathbb{C}}$. Then
(5.6) $\mathcal{D}_{0}\subset\mathcal{D}_{1}\subset\mathcal{D}_{2}\subset\cdots\mathcal{D}_{k}\subset\cdots$
is a filtration on $\mathcal{D}$ such that $\bigcup_{k=0}^{\infty}\mathcal{D}_{k}=\mathcal{D}$ and $\mathcal{D}_{k}\mathcal{D}_{l}\subseteq\mathcal{D}_{k+l}$ for all $k,l\in{\mathbb{Z}}_{+}$. Just as (5.1), we call (5.6) the canonical filtration on $\mathcal{D}$.
Let ${\rm gr(\mathcal{D})}=\bigoplus_{k=0}^{\infty}\mathcal{D}_{k}/\mathcal{D}_{k-1}$ (where $\mathcal{D}_{-1}:=0$), called the associated graded algebra of $\mathcal{D}$ with respect to the canonical filtration, and let $\overline{\mathcal{D}}$ denote the polynomial superalgebra generated by the variables $x^{a}_{i}$, $y^{a}_{r}$, $p_{x^{a}_{i}}$ and $p_{y^{a}_{r}}$, for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$, where $x^{a}_{i}$ and $p_{x^{a}_{i}}$ (resp., $y^{a}_{r}$ and $p_{y^{a}_{r}}$) are even for $1\leq i\leq m$ (resp., $1\leq r\leq p$) and are odd otherwise. It is well known that there is a superalgebra isomorphism
(5.7) ${\rm gr}\!\left(\mathcal{D}\right)\cong\overline{\mathcal{D}}$
sending the elements $x^{a}_{i}+\mathcal{D}_{0}$, $y^{a}_{r}+\mathcal{D}_{0}$, $\partial_{x^{a}_{i}}+\mathcal{D}_{0}$ and $\partial_{y^{a}_{r}}+\mathcal{D}_{0}$ in $\mathcal{D}_{1}/\mathcal{D}_{0}$, respectively, to the elements $x^{a}_{i}$, $y^{a}_{r}$, $p_{x^{a}_{i}}$ and $p_{y^{a}_{r}}$ in $\overline{\mathcal{D}}$, for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$ (see, for example, [CG, Example 1.3.3] and the subsequent discussion). Moreover, the superalgebra ${\rm gr}\!\left(\mathcal{D}\right)$ is naturally a Poisson superalgebra (which is seen as in the case of ${\rm gr}\!\left(U({\mathfrak{g}})\right)$ in Section 5.1), and thus we have a Poisson superalgebra structure on $\overline{\mathcal{D}}$ via (5.7). More explicitly, $\overline{\mathcal{D}}$ is a Poisson superalgebra with the Poisson bracket $\{\cdot,\cdot\}$ given by
(5.8) $\{x^{a}_{i},x^{b}_{j}\}=\{x^{a}_{i},y^{b}_{s}\}=\{y^{a}_{r},y^{b}_{s}\}=\{p_{x^{a}_{i}},p_{x^{b}_{j}}\}=\{p_{x^{a}_{i}},p_{y^{b}_{s}}\}=\{p_{y^{a}_{r}},p_{y^{b}_{s}}\}=0,$
(5.9) $\{p_{x^{a}_{i}},x^{b}_{j}\}=\delta_{i,j}\delta_{a,b},\quad\{p_{x^{a}_{i}},y^{b}_{s}\}=\{p_{y^{a}_{r}},x^{b}_{j}\}=0,\quad\{p_{y^{a}_{r}},y^{b}_{s}\}=\delta_{r,s}\delta_{a,b},$
for $1\leq i,j\leq m+n$, $1\leq r,s\leq p+q$ and $1\leq a,b\leq d$.
The following proposition follows by descending the maps $\phi$ and $\varphi$, given in Proposition 4.1 and Proposition 4.2, to the maps on the associated graded algebras since $\phi(U_{i}({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})))\subseteq\mathcal{D}_{2i}$ and $\varphi(U_{i}({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})))\subseteq\mathcal{D}_{2i}$ for all $i\in{\mathbb{Z}}_{+}$.
Proposition 5.2.
  1. (i)
    The superalgebra homomorphism $\phi:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)\longrightarrow\mathcal{D}$, given in Proposition 4.1, descends to a Poisson superalgebra homomorphism $\overline{\phi}:\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)\longrightarrow\overline{\mathcal{D}}$.
  2. (ii)
    The superalgebra homomorphism $\varphi:U\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)\longrightarrow\mathcal{D}$, given in Proposition 4.2, descends to a Poisson superalgebra homomorphism $\overline{\varphi}:\mathcal{S}\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)\longrightarrow\overline{\mathcal{D}}$.
The maps $\overline{\phi}$ and $\overline{\varphi}$ extend naturally to the Poisson superalgebra homomorphisms
$\overline{\phi}:\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},w^{-1})\!)\longrightarrow\overline{\mathcal{D}}(\!(z^{-1},w^{-1})\!),$
and
$\overline{\varphi}:\mathcal{S}\big({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})\big)(\!(z^{-1},w^{-1})\!)\longrightarrow\overline{\mathcal{D}}(\!(z^{-1},w^{-1})\!),$
respectively. Let $\mu_{{\boldsymbol{\xi}}}^{{\bf w}}\in{\mathfrak{gl}}_{d}^{*}$ and $\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}\in{\mathfrak{gl}}_{\diamondsuit}^{*}$ be as in Section 4.3. We write
$\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\overline{\mathcal{A}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\overline{\mathcal{A}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\overline{\mathcal{A}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$
We also write
$\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\overline{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\overline{{\mathcal{L}}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\overline{{\mathcal{L}}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$
The following theorem gives a duality between the classical Gaudin algebras $\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\overline{\mathcal{A}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$, which we call the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$ for classical Gaudin models with irregular singularities.
Theorem 5.3.
Recall $\gamma_{i}]$, $1\leq i\leq{\boldsymbol{\ell}}$, defined in (4.16). We have
$\prod_{i=1}^{{\boldsymbol{\ell}}}(z-z_{i})^{[\gamma_{i}]}\cdot\overline{\phi}\!\left(\det\!\left(\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)\right)=\prod_{a=1}^{{d^{\prime}}}(w-w_{a})^{\xi_{a}}\cdot\overline{\varphi}\!\left({\rm Ber}^{\bf s}\,\!\left(\overline{{\mathcal{L}}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)\right).$
Consequently, $\overline{\phi}\big(\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\overline{\varphi}\big(\overline{\mathcal{A}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big).$
Proof.
By (5.3),
$\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\bigoplus_{a=1}^{{d^{\prime}}}\big(\!-{\rm J}_{\xi_{a}}(w_{a}-w)\big)-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{e_{ab}\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}}\!\right]_{\!a,b=1,\ldots,d}$
and
$\overline{{\mathcal{L}}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\bigoplus_{i=1}^{{\boldsymbol{\ell}}}\big(\!-{\rm J}_{\gamma_{i}}(z_{i}-z)\big)-\left[(-1)^{|i|}\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}}{(w-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}.$
By the same argument as in Theorem 4.4, we see that
$\prod_{i=1}^{{\boldsymbol{\ell}}}(z-z_{i})^{[\gamma_{i}]}\cdot\overline{\phi}\!\left(\det\!\left(\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})^{t}\right)\right)=\prod_{a=1}^{{d^{\prime}}}(w-w_{a})^{\xi_{a}}\cdot\overline{\varphi}\!\left({\rm Ber}^{\bf s}\,\!\left(\overline{{\mathcal{L}}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)\right)\!.$
Since $\det\!\left(\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})^{t}\right)=\det\!\left(\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)$, the theorem follows. ∎
If any of $p,q,m,n$ is set to 0, we will obtain a special version of the duality. Taking $p=q=n=0$, the identity in Theorem 5.3 recovers [VY, Theorem 3.2]. Taking $p=q=m=0$, it recovers [VY, Theorem 3.4] by Proposition 2.6. In other words, Theorem 5.3 yields the following corollaries.
Corollary 5.4 ([VY, Theorem 3.2]).
Suppose $p=q=n=0$. Then $\overline{\phi}\big(\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\overline{\varphi}\big(\overline{\mathcal{A}}_{m}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big)$.
Corollary 5.5 ([VY, Theorem 3.4]).
Suppose $p=q=m=0$. Then $\overline{\phi}\big(\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\overline{\varphi}\big(\overline{\mathcal{A}}_{0|n}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big)$.
Corollary 5.6.
Set $q=n=0$. Then $\overline{\phi}\big(\overline{\mathcal{A}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\big)=\overline{\varphi}\big(\overline{\mathcal{A}}_{p+m}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\big)$.
Acknowledgments. The first author was partially supported by NSTC grant 113-2115-M-006-010. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2. The authors would like to thank the referee for valuable comments and suggestions.